### v1

# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

Abstract
Foundation models rely on large-scale web-crawled datasets, which frequently contain noisy data, biases, and irrelevant content. Existing data selection techniques typically use human heuristics, downstream evaluation datasets, or specialized scoring models, and can overlook samples’ utility in the training process. Instead, we propose a new approach, Mimic Score, a data quality metric that uses a pretrained reference model as a guide to assess the usefulness of data samples for training a new model. It relies on the alignment between the gradient of the new model parameters and the vector pointing toward the reference model in weight space. Samples that misalign with this direction are considered low-value and can be filtered out. Motivated by the Mimic score, we develop Grad-Mimic, a data selection framework that identifies and prioritizes useful samples, automating the selection process to create effective filters. Empirically, using Mimic scores to guide model training results in consistent performance gains across six image datasets and enhances the performance of CLIP models. Moreover, Mimic scores and their associated filters improve upon existing filtering methods and offer accurate estimation of dataset quality.
1 Introduction
Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], LLaMa-3.1 [3], and GPT-4 [4]. These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content from their web sources. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the model development pipeline [5, 6]. For example, the FineWeb dataset [7], containing 15 trillion tokens used for training large language models, is created through eight carefully designed filtering steps meant to refine raw web content, ensuring high model performance.
While effective, such filtering recipes face several limitations. The choices of filtering steps and rules often rely on handcrafted heuristics and require domain expertise and expensive experimentation. Such filters often fail to provide fine-grained insights into individual samples and overlook samples’ utility in training. These lead to coarse selection and suboptimal model performance.
Existing techniques often suggest selecting samples based on semantic similarity to downstream evaluation datasets [8, 9], using specialized filtering networks [10, 11, 12, 13], or training influence models to score samples [14, 15]. Yet, these approaches require access to additional datasets or involve specialized model training, adding dependencies and complexity.
We introduce the Mimic Score, a new data quality metric that assesses sample contribution in the weight updates using a pretrained reference model. We show that by leveraging a model located in a more optimal part of the weight space, we bypass the need for downstream datasets or specialized training, and can use it as a selection guide. Our approach is based on the alignment between each sample’s gradient and the direction towards reference model in weight space.
Samples that potentially pull the model in undesirable directions—misguiding weight updates—are considered low-value and should be filtered out.
Building on the mimic score, we develop Grad-Mimic, a data selection framework that operates in two stages:
  • Stage 1: During training, Grad-Mimic prioritizes samples to learn by using mimic score, steering the model toward the reference model’s weight space.
  • Stage 2: After training, Grad-Mimic identifies sample utility across training steps and aggregates these assessments for an ensemble filter, automating data selection.
We validate the effectiveness of Grad-Mimic through various experimental setups, presenting empirical results at each stage. First, we create a synthetic and controlled scenario that adds different levels of label noise. We demonstrate its capability to accurately identify mislabeled samples and provide a reliable estimation of overall dataset quality. Compared to the standard training method, Grad-Mimic successfully down-weights the contribution of undesirable samples and enhances model performance across six image datasets.
Next, we test Grad-Mimic in more challenging settings: using large-scale web-crawled multimodal data. We train CLIP models [2] from scratch on 10 million and 100 million samples from the DataComp dataset [16]. We use publicly available CLIP model weights as our reference and demonstrate that mimic scores help navigate training, resulting in performance gains across both scales. Compared to human-designed filters, our refined dataset achieves higher model performance. Furthermore, mimic scores complement CLIP score-based filters [17] by removing low-value samples, lifting model performance with less data.
We summarize our contributions as follows:
  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between per-sample gradient and the vector toward the reference model’s weights.
  • Automated Data Selection Framework: Building on the mimic score, we present Grad-Mimic, a framework that effectively identifies useful samples to improve training and automates effective data selection.
  • Dataset Quality Assessment: The mimic score, along with its developed filters, improves existing filtering strategies and provides reliable estimations of dataset quality.
2 Related Work
Figure 1: High-value vs. Low-value Samples Identified by Mimic Score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their mimic scores. Below each image, we show their caption and corresponding CLIP score (where a higher score suggests good quality [17]). Mimic scores closely align with their CLIP scores. High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
Data Selection.
Data selection techniques can generally be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing the mixture of data domains to curate high-quality datasets [18, 19, 20, 21]. Sample-level methods, which are the focus of Grad-Mimic, aim to filter out individual noisy and unhelpful samples. Prior research has explored various strategies, including identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training update [24, 25]. These methods often rely on batch-level gradients, which can struggle with large amounts of noise in a mini-batch. Grad-Mimic addresses this challenge with the help of a reference model, acting as a reliable guide to select samples.
Data Curation for Multimodal Models.
Multimodal models are typically trained on large-scale web datasets [26, 27, 16, 28], which are often noisy and require careful curation. Previous approaches have included selecting samples based on their semantic similarity to the downstream evaluation datasets [8, 9], using semantic deduplication [29], developing specialized filtering networks [10, 11, 12, 13], or training influence models [14, 15]. While effective, they require access to target datasets or introduce additional training complexities. Grad-Mimic overcomes these limitations and offers a more efficient alternative: using pretrained weights to identify useful samples.
Weak Supervision.
Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [30, 31, 32, 33, 34]. These estimates typically come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [35, 36], often encoded as labeling functions. The outputs of these labeling functions are modeled to assess their reliability and then aggregated to produce final labels [30]. Weak supervision has demonstrated success in various domains, including segmentation tasks [37], relation extraction [38], MRI sequence analysis [39], sensor data [40], chemical reaction extraction [41], and enhancing predictions from pretrained models [42, 43]. Unlike most existing works focusing on label aggregation for data annotation, Grad-Mimic uses weak supervision for filtering purposes. We mitigate assessment noise across training steps, aggregating filter outputs into an effective ensemble filter.
3 Evaluating Sample Utility
Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients. Then it adapts alignment scores to reweight gradients that contributed to the weight movements. In Stage 2, derived alignment scores, mimic scores, are used to identify low-value samples and build an ensemble filter, enabling accurate noise detection, precise dataset quality assessment, and training dataset discovery.
We quantify samples by their contributions to the learning process. Our principle is that samples that potentially pull the model in undesirable directions, thereby misdirecting weight updates, should be considered low-value.
We start with setup and notation, then explain how our scoring metric is derived and used in our data selection framework, Grad-Mimic.
Notation.
Let $D=\{s_{i}\}_{i=1}^{n}$ denote a dataset of $n$ samples drawn from a distribution $\mathcal{D}$ supported on the space $\mathcal{S}$. At training step $t$, model parameters $\theta_{t}$, are iteratively optimized using the dataset $D$. While our framework supports various training settings, we focus on supervised learning for clarity. We assume $\mathcal{S}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. Each sample $s_{i}$ can be expressed as $(x_{i},y_{i})$, where noise may be present either in the instance $x_{i}$ or in the label $y_{i}$. The empirical loss across samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$. A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
3.1 Mimic Score Calculation
To evaluate whether a sample steers the model in an undesirable direction, we use a pretrained model as a reference. This reference model, denoted by $\theta_{\text{ref}}$, resides in a more optima part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample’s utility in approximating a better weight configuration. These reference weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [44, 45].
The reference model can be obtained in two ways: either by training it on the dataset to achieve a target performance, if resources permit, or more efficiently, by using publicly available pretrained models, which eliminate the need for data access and expensive training.
The vector pointing toward the reference model, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$. We examine how each sample’s negative gradient $-g_{i,t}$, intended for updating model weights, aligns with vector $v_{t}$. We measure the alignment degree by considering both the direction and magnitude of the negative gradient. Specifically, we compute the projection length of $-g_{i,t}$ onto $v_{t}$, yielding an alignment score $m_{i,t}$, computed as follows
[…]
This alignment score, named the mimic score, reflects how much a sample can drive the model closer to the reference model. A sample having a lower mimic score suggests it has limited utility in guiding model updates, making it a potential candidate for exclusion from future training.
4 Grad-Mimic: Data Selection Framework
The mimic score represents each sample’s usefulness in guiding the model in a better direction. Building on this metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes samples to learn and then identifies useful samples to design an effective filter. A complete Grad-Mimic workflow is illustrated in Fig. 2.
4.1 Stage 1: Reweighted Gradient Updates
We first use mimic scores to aid model training by re-weighting sample gradients. Unlike standard gradient descent, which assigns equal weight to all the samples in the mini-batch, Grad-Mimic uses mimic scores to amplify helpful samples and down-weight unhelpful ones.
To achieve this, each sample’s mimic score is first normalized using the softmax function with a temperature parameter, $\tau$. The normalized score for a sample $s_{i}$ in a batch of size $b$ is computed as
$\overline{m}_{i,t}=\frac{e^{-m_{i,t}/\tau}}{\sum_{j=1}^{b}e^{-m_{j,t}/\tau}}.$ (2)
Then, the weight update step in Grad-Mimic is modified as
$\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}.$ (3)
The temperature $\tau$ controls the sensitivity of sample reweighting, allowing us to adjust how sharply the model prioritizes samples. A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher temperature encourages Grad-Mimic converges to standard gradient descent.
4.2 Stage 2: Automated Filter Design
The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, automate data selection based on these assessments, and estimate dataset quality.
4.2.1 Sample Identification
Grad-Mimic first gathers normalized mimic scores for each sample at every training step. These scores indicate the sample contributions over training, which allows us to decide whether to retain or discard a sample. Grad-Mimic supports several sample identification methods:
  • Threshold-based Selection: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
  • 1D Clustering: Samples are categorized into two groups using clustering techniques such as $k$-means clustering [46] or Gaussian Mixture Models (GMMs), allowing more automated selection based on their assigned cluster.
  • Top-k Percent Selection: Samples are ranked by mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
After identification, we binarize mimic scores and categorize samples into “retain” or “discard” at each training step. We treat these assigned groups as filter outputs and then combine them for a final filter.
4.2.2 Filter Output Aggregation
While filtering at an individual training step can refine the dataset, relying on isolated filters risks overlooking training dynamics. Additionally, considering sample utility across different stages may yield complementary signals.
To address this, Grad-Mimic leverages weak supervision techniques [30, 31, 32, 33, 34], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps. We begin with learning a model to evaluate the reliability of each step assessment [32]. Once established, this model aggregates filtering decisions into a high-quality ensemble filter. Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics. Grad-Mimic employs the Snorkel framework [31], a widely used method in the weak supervision community.
Ultimately, the ensemble filter selects high-value samples for a refined dataset. We can estimate dataset quality by analyzing the proportion of retained samples or using samples’ mimic scores as an overall metric. We summarize steps in Grad-Mimic in Algorithm 1 placed in Appendix B.
5 Experiments
We assess the effectiveness of mimic score within Grad-Mimic using two experimental setups across datasets of varying scales and domains. We first test in a controllable setting, adding noise into sample labels (Sec. 5.1), followed by evaluations on large-scale web-crawled datasets (Sec. 5.2). Our goals are to validate the following claims in both setups:
  • Enhanced Model Performance: The reference model acts as a reliable guide. Training a new model with samples prioritized by mimic scores enhances performance.
  • Accurate Sample Identification: Grad-Mimic effectively detects noisy samples, automates data selection, and complements existing filtering methods.
  • Dataset Quality Assessment: Mimic scores and their associated filters offer reliable estimates of dataset quality, highly correlated with noise levels and performance gains.
5.1 Simulating Mislabeled Samples
Setups.
We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [47], Flowers102 [48], STL10 [49], OxfordIIIT Pet [50], CIFAR10 [51], and CIFAR100 [51]. We fine-tune ViT-B/16 models [52] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last layer weights as the reference to navigate training on the noisy datasets. We detail more training configurations in Appendix C.1.
After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and clustering methods using $k$-means and GMM, then aggregate filter outputs across training steps using Snorkel framework [31].
Expected Results.
We aim to validate that a pretrained reference model can serve as a strong guide for model training. Moreover, we expect that mimic scores obtained from Grad-Mimic can help identify low-value, mislabeled samples.
Full Fine-tuning DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 65.59 91.01 93.20 82.88 94.89 81.34 84.82
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
SGD 46.91 42.02 33.24 65.54 55.90 39.84 75.35 58.57 49.98 70.43 64.62 55.63 90.31 86.69 82.01 69.42 63.77 59.28 69.66 61.93 53.33
GraNd 29.73 8.83 15.43 10.80 2.78 27.92 78.39 52.32 64.30 56.83 45.38 45.43 87.56 83.79 80.04 18.65 8.83 7.84 46.99 33.66 40.16
AGRA 46.28 37.29 32.39 0.62 0.93 0.82 84.35 79.94 78.17 76.97 77.11 69.39 89.23 86.79 68.09 11.96 8.44 6.88 51.57 48.42 42.62
Grad-Match 48.88 42.66 33.24 37.92 54.94 42.62 83.86 81.16 77.18 74.35 61.73 51.16 44.48 35.69 37.59 69.41 66.21 20.09 58.82 57.07 43.65
Grad-Mimic 49.20 42.82 33.83 68.58 56.46 44.27 72.06 71.85 83.09 81.98 78.30 73.34 90.52 89.07 66.41 73.97 74.31 24.02 72.72 68.60 54.16
Linear Probing DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
GraNd 36.22 30.59 25.32 18.23 13.90 10.49 68.70 59.55 49.68 69.47 60.02 48.13 88.02 85.61 81.48 71.60 70.51 69.27 58.71 53.36 47.40
AGRA 41.81 36.28 31.49 41.81 36.28 31.49 96.19 95.15 93.11 85.25 82.53 78.39 92.51 92.01 90.99 72.50 71.30 69.69 71.68 68.93 65.86
Grad-Match 51.81 47.55 41.33 27.00 20.21 14.90 96.06 95.34 93.44 86.86 85.75 83.18 92.62 92.04 91.17 75.39 74.46 73.18 71.62 69.23 66.20
Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, leading to improved denoising and performance gains.
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
Threshold ($1/32$) 85.13 88.91 91.82 90.93 94.9 96.92 90.31 94.11 97.21 94.41 96.86 97.93 98.37 98.19 97.19 97.66 97.57 95.80
1D $k$-means 77.13 76.68 77.17 84.45 87.62 87.47 78.90 83.04 78.82 95.80 95.49 93.73 98.62 98.17 97.56 98.10 97.46 96.62
GMM 97.85 96.72 95.68 98.39 96.96 94.21 97.96 97.16 95.69 98.04 97.07 95.86 95.70 92.89 88.62 95.86 94.31 92.55
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric. Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and identification methods. Notably, its performance remains high and robust regardless of the noise level.
Figure 3: Mimic Score Distribution: Extracted from CIFAR100 dataset with a noise level of 0.5. The distributions clearly separate, aligning with the correctness of the labels.
Baselines.
We compare Grad-Mimic’s training method with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of a random batch, then excludes outlier samples. (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch.
Stage 1 Results.
We present Grad-Mimic’s two-stage results separately. The Stage 1 results, shown in Table 1, report testing accuracy across different methods and noisy datasets using full fine-tuning and linear probing. We evaluate Grad-Mimic’s robustness by averaging accuracy across datasets under different levels of label noise. Results show that Grad-Mimic effectively identifies and down-weights mislabeled samples, improving denoising and closing the performance gap to the reference model. Notably, in the full fine-tuning setup where we only target the last layer to mimic, we successfully guide the entire model weight updates, ultimately outperforming other methods. This also highlights the computational feasibility of our approach. These findings validate the idea that using the direction toward a better-performing model can guide model training and lead to enhanced performance.
Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
ImageNet
dist. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
Standard Training 0.026 0.035 0.139 0.114 0.131
0.03 0.026 0.035 0.153 0.115 0.136
0.05 0.026 0.035 0.147 0.114 0.135
0.07 0.026 0.035 0.151 0.116 0.135
0.3 0.026 0.034 0.154 0.112 0.137
Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
0.03 0.025 0.035 0.149 0.114 0.133
0.05 0.026 0.037 0.162 0.118 0.146
0.07 0.027 0.036 0.166 0.118 0.145
0.3 0.026 0.035 0.161 0.114 0.144
Small Grad-Mimic Last MLP Layer in Image Encoder 0.5 0.027 0.035 0.160 0.117 0.145
Standard Training 0.171 0.148 0.253 0.217 0.254
Medium Grad-Mimic
Last MLP Layer
in Image Encoder
0.05 0.169 0.151 0.262 0.216 0.258
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance.
((a))
((b))
Figure 4: Mimic scores offer precise dataset quality estimation: Left figure (a) computes the portion of retained samples to evaluate dataset quality in simulation experiments. Right figure (b) takes the average on mimic scores in each filtered dataset for estimation. [S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
Stage 2 Results.
Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled samples across all datasets and demonstrates robust performance when facing different noise levels.
We use CIFAR100 dataset as an example and visualize mimic score distribution at each training step in Figure 3. In this dataset, half of the labels are flipped. We see their distributions clearly separate into two groups, confirming that mimic scores can serve as informative signals to identify samples. Despite variations over training, these dynamics can be effectively captured in our aggregation step.
We use an ensemble filter from each dataset created by GMM clustering to estimate dataset quality using the proportion of retained data. The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903. This demonstrates mimic score’s effectiveness for dataset quality estimation.
5.2 Data Selection in Large-scale Web Datasets
Setups.
We evaluate Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small- and medium-scale DataComp datasets [16], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, for CLIP model pretraining [2]. We follow the training setup from DataComp [53, 16] and use publicly available pretrained CLIP weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This pretrained model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy to the ideal reference point for scenarios where training such powerful models is infeasible. We target to mimic the final MLP layer in the text and image encoders respectively. We evaluate model performance using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Additional training details can be found in Appendix C.2.
In Stage 2, we use top-$k$ percent selection method to create the final filter, subsequently training a CLIP model on the curated dataset to validate filter’s effectiveness.
Expected Results.
We anticipate mimic scores help large-scale training focus on high-value samples, while the derived scores will be instrumental in automating effective data curation.
Baselines.
In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [16]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [17]: selecting top-$k$ percent samples based on embedding alignment between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2]. We see CLIP score as a strong baseline, which is prebuilt through heavy training using hundreds of millions of samples.
Stage 1 Results.
Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training across all the temperature settings. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. These results highlight Grad-Mimic’s success in using pretrained model weights as a guide to prioritize high-value samples to learn.
Scale Filtering Strategy Dataset Size ImageNet
ImageNet
dis. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
No Filtering 10.7M 0.026 0.035 0.139 0.114 0.131
Basic Filtering 3M 0.031 0.040 0.159 0.115 0.141
Mimic Score Top 30% 3M 0.028 0.037 0.151 0.107 0.135
Mimic Score Top 35% 3.8M 0.026 0.037 0.158 0.107 0.143
Mimic Score Top 40% 4.2M 0.028 0.035 0.151 0.110 0.139
CLIP Score (B/32 Top 30%) 3M 0.051 0.054 0.183 0.118 0.165
CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M 0.047 0.053 0.174 0.106 0.163
Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M 0.047 0.052 0.173 0.113 0.164
No Filtering 101.9M 0.026 0.035 0.139 0.114 0.254
Basic Filtering 30M 0.209 0.176 0.264 0.232 0.267
Mimic Score Top 30% 29M 0.192 0.163 0.269 0.218 0.268
CLIP Score (B/32 Top 30%) 30M 0.275 0.229 0.332 0.246 0.321
CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M 0.270 0.227 0.339 0.246 0.323
CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M 0.274 0.226 0.334 0.243 0.323
Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M 0.264 0.223 0.330 0.244 0.322
Table 4: Stage 2 Results in DataComp Experiment: Mimic score-based filters perform better than basic filtering and complement CLIP score-based filters by removing low-value samples. Symbol $\setminus$” denotes the exclusion of curated datasets.
Stage 2 Results.
We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic successfully automates the selection process, outperforming the basic filtering strategy at both dataset scales. Additionally, we remove less useful samples identified by Grad-Mimic to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. Besides, we include the top-ranked approach during our time of experimentation [54], which uses an orthogonal approach based on object detection models. By calculating the average mimic score in each curated dataset, we compare with their corresponding performance gains. The results, shown in Figure 4, reveal a strong alignment between performance gains and average mimic scores, with a Pearson correlation of 0.958, demonstrating that mimic scores can serve as a reliable metric for estimating dataset quality.
Figure 5: The filtering logic behind mimic score aligns with handcrafted heuristics, particularly in caption length. Low-value samples (in the left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
((a))
((b))
Figure 6: Mimic score distribution and its relationship with CLIP score.
Sample Analysis.
We analyze high- and low-value samples ranked by mimic score. In Figure 5, we compare the token usage distribution in captions. We find that bottom-ranked samples often carry captions that are either too short (one or two words) or excessively long (exceeding the maximum token limit in CLIP’s tokenizer), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or visually incoherent images in the medium-scale dataset (see examples shown in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic framework.
We select the top 30% and bottom 30% of samples ranked by their mimic scores and visualize their distribution along with corresponding CLIP scores. In Figure 6, we observe high-value samples distributed towards higher CLIP scores, while low-value samples are more concentrated in lower CLIP score ranges (half of the samples are scored less than 0.2). Furthermore, we calculate their correlation using randomly selected samples. For a better visualization, we normalize both scores into the range 0 to 1. We see a positive correlation between them, supporting the reliability of mimic scores in identifying high-quality samples.
Filtering Strategy Random Selection Mimic Score Selection
Jaccard
Similarity
Percentage of
Overlap
Jaccard
Similarity
Percentage of
Overlap
CLIP-ViT B/32 Top 30% 0.149 0.258 0.232 0.376
CLIP-ViT L/14 Top 30% 0.150 0.260 0.236 0.382
DataComp’23 Top-ranked 0.166 0.284 0.271 0.426
Table 5: Mimic score can help in predicting whether a sample was used to train the reference model.
Predicting Pretraining Dataset
We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we use their final layer weights as the reference and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic scores appear in their curated dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking the final layer weights without direct access to their filtering steps.
6 Conclusion
We introduce a new data quality metric called Mimic Score, designed to assess samples’ utility in the training process. We quantify samples by measuring the alignment of per-sample gradient and the vector toward a pretrained reference model. Building on Mimic Score, we develop Grad-Mimic to automate data selection process. Empirically, we demonstrate that Grad-Mimic effectively prioritizes identified samples to learn, resulting in enhanced model performance. Moreover, Mimic Score can serve as a reliable metric for creating effective filters and estimating dataset quality. Developed filters outperform human-designed ones and exhibit a strong correlation with performance gains.
7 Acknowledgments
We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, and John Cooper for their helpful feedback and valuable discussion.
Appendix A Discussion
A.1 Broader Impacts and Limitations.
The importance of data selection has grown in building foundation models. We demonstrate that Grad-Mimic can effectively identify and exclude noisy or unhelpful samples. This leads to reduced computational and resource costs. Besides, Grad-Mimic provides the second benefit by enabling models to focus their training on high-value samples, which contributes to performance improvements.
We do not foresee explicit negative impacts arising from Grad-Mimic. However, if the reference model used for sample selection is unreliable, there remains a risk of refining suboptimal datasets. To address this, incorporating human expert-designed filtering rules may be necessary to ensure the quality of curated datasets.
Appendix B Grad-Mimic Algorithm
We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights. Our algorithm supports various training settings, such as supervised learning and self-supervised learning. We have validated both training scenarios through simulation experiments (Sec. 5.1) and by training on web-crawled datasets with CLIP models (Sec.5.2). Moreover, these experiments evaluate Grad-Mimic performance across three different training configurations: full-parameter fine-tuning, linear probing, and training models from scratch.
Algorithm 1 Grad-Mimic (Stage 1)
reference model $\theta_{\text{ref}}$, new model $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature $\tau$
samples’ mimic scores $\overline{m}_{i,t}$, trained model $\theta_{T}$
Initialize new weights $\theta_{0}$
for $t\in[T]$ do
     uniformly sample a batch $\{s_{i}\}_{i=1}^{b}$ from $D$
     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as guide
     for $i\in[b]$ do
         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradients
         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|_{2}}$triangleright compute mimic scores
     end for
     $\overline{m}_{:,t}=\frac{e^{-m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{-m_{j,t}/\tau}}$triangleright normalize mimic scores
     $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright reweight gradients
end for
Appendix C Experimental Details
We provide more details about our training setups, and computational resources.
C.1 Simulation Experiments
In the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [55]. Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs. We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper. Results for other temperature values are presented in Table 6. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets using identical configurations but with different random seeds for weight initialization. These experiments were performed on a Nvidia Tesla A100.
Temperature $\tau$ DTD Flower102 Oxford-IIIT Pet CIFAR10 CIFAR100
SGD 52.8 35.6 87.9 93.2 76.2
Grad-Match 53.1 32.1 87.4 93.0 76.1
1.0 56.5 44.9 89.0 94.1 77.0
0.9 56.4 45.3 89.1 94.1 77.1
0.8 56.3 45.8 89.1 94.1 77.2
0.7 56.3 46.1 89.2 94.2 77.2
Grad-Mimic 0.6 56.0 46.4 89.3 94.2 77.3
Table 6: Stage 1 Results in Simulation Experiment with Different Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
C.2 DataComp Experiments
We adopt the training setup from DataComp [53, 16]. CLIP models are trained from scratch using a contrastive objective over image-caption pairs. Each model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset. These experiments ran on 8 Nvidia Tesla A100s.
We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders. The layers used are “visual.transformer.resblocks.11.mlp.c_fc.weight” and “text.transformer.resblocks.11.mlp.c_fc.weight”. Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [16] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5). Results are presented in Table 3.
Appendix D Ablation Studies
The Choice of Temperature.
We present Grad-Mimic results with different temperature values and compare them to baseline methods (SGD and Grad-Match [24]). The results in the simulation experiments are presented in Table 6. We set the noise level to 0.3 and fine-tune ViT model under linear probing configuration. Grad-Mimic outperforms the baseline methods across all temperature settings. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.

### v2

@@
 Abstract
-Foundation models rely on large-scale web-crawled datasets, which frequently contain noisy data, biases, and irrelevant content.
-Existing data selection techniques typically use human heuristics, downstream evaluation datasets, or specialized scoring models, and can overlook samples’ utility in the training process.
-Instead, we propose a new approach, Mimic Score, a data quality metric that uses a pretrained reference model as a guide to assess the usefulness of data samples for training a new model.
-It relies on the alignment between the gradient of the new model parameters and the vector pointing toward the reference model in weight space.
-Samples that misalign with this direction are considered low-value and can be filtered out.
-Motivated by the Mimic score, we develop Grad-Mimic, a data selection framework that identifies and prioritizes useful samples, automating the selection process to create effective filters.
-Empirically, using Mimic scores to guide model training results in consistent performance gains across six image datasets and enhances the performance of CLIP models.
-Moreover, Mimic scores and their associated filters improve upon existing filtering methods and offer accurate estimation of dataset quality.
+Foundation models are trained on large-scale web-crawled datasets, which often contain noise, biases, and irrelevant information.
+This motivates the use of data selection techniques, which can be divided into model-free variants—relying on heuristic rules and downstream datasets—and model-based, e.g., using influence functions.
+The former can be expensive to design and risk introducing unwanted dependencies, while the latter are often computationally prohibitive.
+Instead, we propose an efficient, model-based approach using the Mimic Score, a new data quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model.
+It relies on the alignment between gradients and a target direction induced by the reference model.
+Using the derived Mimic Scores, we develop Grad-Mimic, a framework that prioritizes samples for learning, creates effective filters, and automates data selection.
+Empirically, using Mimic Scores to guide training improves data efficiency, results in consistent performance gains across six image datasets, and includes enhancements to CLIP models.
+Moreover, Mimic Score-based filters improve upon existing filtering methods, e.g., cutting 4.7 million samples to train better CLIP models while offering accurate estimation of training dataset quality.
 1 Introduction
-Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], LLaMa-3.1 [3], and GPT-4 [4].
-These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content from their web sources.
-To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the model development pipeline [5, 6].
-For example, the FineWeb dataset [7], containing 15 trillion tokens used for training large language models, is created through eight carefully designed filtering steps meant to refine raw web content, ensuring high model performance.
-While effective, such filtering recipes face several limitations.
-The choices of filtering steps and rules often rely on handcrafted heuristics and require domain expertise and expensive experimentation.
-Such filters often fail to provide fine-grained insights into individual samples and overlook samples’ utility in training.
-These lead to coarse selection and suboptimal model performance.
-Existing techniques often suggest selecting samples based on semantic similarity to downstream evaluation datasets [8, 9], using specialized filtering networks [10, 11, 12, 13], or training influence models to score samples [14, 15].
-Yet, these approaches require access to additional datasets or involve specialized model training, adding dependencies and complexity.
-We introduce the Mimic Score, a new data quality metric that assesses sample contribution in the weight updates using a pretrained reference model.
-We show that by leveraging a model located in a more optimal part of the weight space, we bypass the need for downstream datasets or specialized training, and can use it as a selection guide.
-Our approach is based on the alignment between each sample’s gradient and the direction towards reference model in weight space.
-Samples that potentially pull the model in undesirable directions—misguiding weight updates—are considered low-value and should be filtered out.
+Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], and GPT-4 [3].
+These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content.
+To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [4, 5].
+For example, the FineWeb dataset [6], with 15 trillion tokens, is refined through eight carefully designed filtering steps to ensure high model performance.
+Existing data selection methods have several limitations.
+We can divide these into two categories: those that are independent of a model being trained (or using pretrained models) and those that rely on such a model.
+Model-free approaches, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [7, 8] often require expensive experimentation to establish selection rules and add unwanted dependencies on additional datasets.
+Furthermore, these methods can fail to provide fine-grained insights into individual samples and overlook sample utility in training, resulting in suboptimal performance.
+In contrast, model-based techniques, e.g., training specialized filtering networks  [9, 10, 11, 12] or using data influence functions  [13, 14], introduce complexity and are often computationally expensive.
+Figure 1: High-value vs.
+low-value samples identified by the Mimic score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their Mimic scores.
+Each image is accompanied by its caption and corresponding CLIP score (higher scores suggest better quality).
+Mimic scores align closely with CLIP scores.
+High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
+To address these limitations, we introduce Mimic Score, a new data quality metric that assesses sample’s contribution to weight updates using reference model’s weights.
+This is an efficient model-based technique.
+We show that by leveraging pretrained model weights (located in a more optimal part of the weight space), we bypass the need for downstream datasets or specialized model training and can use it as our selection guide.
+Our approach is based on the alignment between each sample’s negative gradient and a direction pointing toward the reference model in the weight space.
+We argue that samples that potentially pull the model in undesirable directions—thereby misguiding weight updates—should be considered low-value and ruled out.
+Leveraging model weights to evaluate sample utility offers several advantages.
+First, with the growing availability of well-trained models, adopting model weights has become more accessible than ever.
+Second, unlike other reference model-based approaches that require loading a second model and performing full inference passes [15, 16], our method scores each sample directly from the weights, improving efficiency.
 Building on the mimic score, we develop Grad-Mimic, a data selection framework that operates in two stages:
-  • Stage 1: During training, Grad-Mimic prioritizes samples to learn by using mimic score, steering the model toward the reference model’s weight space.
-  • Stage 2: After training, Grad-Mimic identifies sample utility across training steps and aggregates these assessments for an ensemble filter, automating data selection.
-We validate the effectiveness of Grad-Mimic through various experimental setups, presenting empirical results at each stage.
-First, we create a synthetic and controlled scenario that adds different levels of label noise.
-We demonstrate its capability to accurately identify mislabeled samples and provide a reliable estimation of overall dataset quality.
-Compared to the standard training method, Grad-Mimic successfully down-weights the contribution of undesirable samples and enhances model performance across six image datasets.
-Next, we test Grad-Mimic in more challenging settings: using large-scale web-crawled multimodal data.
-We train CLIP models [2] from scratch on 10 million and 100 million samples from the DataComp dataset [16].
-We use publicly available CLIP model weights as our reference and demonstrate that mimic scores help navigate training, resulting in performance gains across both scales.
-Compared to human-designed filters, our refined dataset achieves higher model performance.
-Furthermore, mimic scores complement CLIP score-based filters [17] by removing low-value samples, lifting model performance with less data.
+  • Stage 1: During training, Grad-Mimic prioritizes which samples to learn based on their mimic scores, improving data efficiency while steering the model toward the reference model’s weight space.
+  • Stage 2: After training, Grad-Mimic uses mimic scores to evaluate each sample’s utility across training steps.
+These assessments are then aggregated to create an ensemble filter, which can further refine the dataset.
+We validate Grad-Mimic through diverse experiments.
+First, in a controlled setting with varying levels of label noise, Grad-Mimic accurately identifies mislabeled samples and estimates training dataset quality.
+During training, it effectively down-weights undesirable samples and improves data efficiency across six image datasets.
+Next, we test it on large-scale web-crawled multimodal data by training CLIP models from scratch [2] on 10M and 100M samples from DataComp  [17].
+Using publicly available pretrained weights as our reference, mimic scores effectively navigate training.
+The resulting filter performs better than human-designed ones and can complement CLIP score-based filters [18], lifting CLIP model performance with less training data.
 We summarize our contributions as follows:
-  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between per-sample gradient and the vector toward the reference model’s weights.
-  • Automated Data Selection Framework: Building on the mimic score, we present Grad-Mimic, a framework that effectively identifies useful samples to improve training and automates effective data selection.
-  • Dataset Quality Assessment: The mimic score, along with its developed filters, improves existing filtering strategies and provides reliable estimations of dataset quality.
+  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between gradients and the vector toward the reference model in the weight space.
+  • Automated Data Selection Framework: Using the mimic score, we present Grad-Mimic, a framework that effectively identifizes useful samples to improve data efficiency and automates effective data selection.
+  • Dataset Quality Assessment: The mimic score, along with its resulting filters, improves existing filtering strategies and provides reliable estimation of dataset quality.
 2 Related Work
-Figure 1: High-value vs.
-Low-value Samples Identified by Mimic Score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their mimic scores.
-Below each image, we show their caption and corresponding CLIP score (where a higher score suggests good quality [17]).
-Mimic scores closely align with their CLIP scores.
-High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
+Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients.
+Then it adapts alignment scores to reweight gradients that contributed to the weight updates.
+In Stage 2, derived alignment scores—mimic scores—are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality estimation, and training dataset discovery.
 Data Selection.
-Data selection techniques can generally be categorized into group- and sample-level approaches.
-Group-level approaches focus on optimizing the mixture of data domains to curate high-quality datasets [18, 19, 20, 21].
-Sample-level methods, which are the focus of Grad-Mimic, aim to filter out individual noisy and unhelpful samples.
-Prior research has explored various strategies, including identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training update [24, 25].
-These methods often rely on batch-level gradients, which can struggle with large amounts of noise in a mini-batch.
-Grad-Mimic addresses this challenge with the help of a reference model, acting as a reliable guide to select samples.
+Data selection techniques can be categorized into group- and sample-level approaches.
+Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [19, 20, 21].
+Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and filter out unhelpful samples.
+Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training [24, 25].
+These methods can struggle with large amounts of noise in batch gradients.
+Grad-Mimic overcome this with the help of a reference model.
 Data Curation for Multimodal Models.
-Multimodal models are typically trained on large-scale web datasets [26, 27, 16, 28], which are often noisy and require careful curation.
-Previous approaches have included selecting samples based on their semantic similarity to the downstream evaluation datasets [8, 9], using semantic deduplication [29], developing specialized filtering networks [10, 11, 12, 13], or training influence models [14, 15].
-While effective, they require access to target datasets or introduce additional training complexities.
-Grad-Mimic overcomes these limitations and offers a more efficient alternative: using pretrained weights to identify useful samples.
+Multimodal models rely on large-scale web datasets [26, 17, 27], which are often naturally noisy and require careful curation.
+Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [7, 8], using semantic deduplication [28], developing specialized filtering networks [9, 10, 11, 12], and using data influence functions [13, 14].
+While effective, they require access to target datasets, introduce training complexities, and can be expensive in computation.
+Grad-Mimic overcomes these and offers a more efficient alternative: using public available model weights to identify useful samples.
 Weak Supervision.
-Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [30, 31, 32, 33, 34].
-These estimates typically come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [35, 36], often encoded as labeling functions.
-The outputs of these labeling functions are modeled to assess their reliability and then aggregated to produce final labels [30].
-Weak supervision has demonstrated success in various domains, including segmentation tasks [37], relation extraction [38], MRI sequence analysis [39], sensor data [40], chemical reaction extraction [41], and enhancing predictions from pretrained models [42, 43].
-Unlike most existing works focusing on label aggregation for data annotation, Grad-Mimic uses weak supervision for filtering purposes.
-We mitigate assessment noise across training steps, aggregating filter outputs into an effective ensemble filter.
+Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [29, 30, 31, 32].
+These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [33, 34], often encoded as labeling functions.
+The outputs of these labeling functions are modeled and then aggregated to produce final labels [29].
+Weak supervision has demonstrated success across diverse domains [35, 36, 37, 38, 39].
+Most prior works focus on weak label aggregation to construct datasets, Grad-Mimic leverages weak supervision for a new purpose: modeling sample utility assessments for a combined filtering decision.
 3 Evaluating Sample Utility
-Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients.
-Then it adapts alignment scores to reweight gradients that contributed to the weight movements.
-In Stage 2, derived alignment scores, mimic scores, are used to identify low-value samples and build an ensemble filter, enabling accurate noise detection, precise dataset quality assessment, and training dataset discovery.
 We quantify samples by their contributions to the learning process.
-Our principle is that samples that potentially pull the model in undesirable directions, thereby misdirecting weight updates, should be considered low-value.
-We start with setup and notation, then explain how our scoring metric is derived and used in our data selection framework, Grad-Mimic.
+Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value.
+We start with setup and notation, then explain how the scoring metric is derived and used in our data selection framework, Grad-Mimic.
 Notation.
@@
 3.1 Mimic Score Calculation
-To evaluate whether a sample steers the model in an undesirable direction, we use a pretrained model as a reference.
-This reference model, denoted by $\theta_{\text{ref}}$, resides in a more optima part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$.
-We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample’s utility in approximating a better weight configuration.
-These reference weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [44, 45].
-The reference model can be obtained in two ways: either by training it on the dataset to achieve a target performance, if resources permit, or more efficiently, by using publicly available pretrained models, which eliminate the need for data access and expensive training.
-The vector pointing toward the reference model, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$.
+To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our guide.
+These reference weights, denoted by $\theta_{\text{ref}}$, reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$.
+We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$).
+These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [40, 41].
+Reference weights can be obtained in two ways: by training a model on the dataset to achieve a desired performance level, if resources permit, or more efficiently, by using publicly available well-trained models, which are now more accessible and can eliminate the need for direct data access or possible costly training.
+The vector pointing toward reference model’s weights, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$.
 We examine how each sample’s negative gradient $-g_{i,t}$, intended for updating model weights, aligns with vector $v_{t}$.
-We measure the alignment degree by considering both the direction and magnitude of the negative gradient.
+We measure this alignment degree by considering both the direction and magnitude of the negative gradient.
 Specifically, we compute the projection length of $-g_{i,t}$ onto $v_{t}$, yielding an alignment score $m_{i,t}$, computed as follows
@@
 4 Grad-Mimic: Data Selection Framework
-The mimic score represents each sample’s usefulness in guiding the model in a better direction.
-Building on this metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes samples to learn and then identifies useful samples to design an effective filter.
-A complete Grad-Mimic workflow is illustrated in Fig.
+Building on this simple but effective metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn and then identifies useful samples to design an effective filter.
+The complete Grad-Mimic workflow is illustrated in Fig.
 2.
@@
 The normalized score for a sample $s_{i}$ in a batch of size $b$ is computed as
-$\overline{m}_{i,t}=\frac{e^{-m_{i,t}/\tau}}{\sum_{j=1}^{b}e^{-m_{j,t}/\tau}}.$ (2)
+$\overline{m}_{i,t}=\frac{e^{m_{i,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}.$ (2)
 Then, the weight update step in Grad-Mimic is modified as
@@
 4.2 Stage 2: Automated Filter Design
-The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, automate data selection based on these assessments, and estimate dataset quality.
+The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, creating a combined filter based on sample utility assessments.
 4.2.1 Sample Identification
@@
 Grad-Mimic supports several sample identification methods:
-  • Threshold-based Selection: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
-  • 1D Clustering: Samples are categorized into two groups using clustering techniques such as $k$-means clustering [46] or Gaussian Mixture Models (GMMs), allowing more automated selection based on their assigned cluster.
-  • Top-k Percent Selection: Samples are ranked by mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
-After identification, we binarize mimic scores and categorize samples into “retain” or “discard” at each training step.
-We treat these assigned groups as filter outputs and then combine them for a final filter.
+  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
+  • 1D Clustering: Samples are categorized into two groups using clustering such as $k$-means clustering [42] or Gaussian Mixture Models (GMMs), allowing more automated identification based on their assigned cluster.
+  • Top-k Percent: Samples are ranked by their mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
+After identification, normalized mimic scores are binarized, and samples are categorized into “retain” or “discard” at each training step.
+We treat these assigned groups as each training step’s votes and then combine for a final filter.
 4.2.2 Filter Output Aggregation
-While filtering at an individual training step can refine the dataset, relying on isolated filters risks overlooking training dynamics.
-Additionally, considering sample utility across different stages may yield complementary signals.
-To address this, Grad-Mimic leverages weak supervision techniques [30, 31, 32, 33, 34], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps.
-We begin with learning a model to evaluate the reliability of each step assessment [32].
+Although using the assessment at a training step can refine the dataset, relying on such isolated filters risks overlooking training dynamics.
+Moreover, evaluating a sample’s utility across all training stages may yield complementary signals.
+To address this, Grad-Mimic uses weak supervision techniques [29, 30, 31, 32], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps.
+We start by learning a model to evaluate the reliability of each step’s assessment [31].
 Once established, this model aggregates filtering decisions into a high-quality ensemble filter.
 Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics.
-Grad-Mimic employs the Snorkel framework [31], a widely used method in the weak supervision community.
-Ultimately, the ensemble filter selects high-value samples for a refined dataset.
-We can estimate dataset quality by analyzing the proportion of retained samples or using samples’ mimic scores as an overall metric.
-We summarize steps in Grad-Mimic in Algorithm 1 placed in Appendix B.
+We employ the Snorkel framework [30], a widely used method in the weak supervision community.
+Ultimately, we can use this ensemble filter to select high-value samples for a refined dataset.
+Moreover, we can use it to estimate training dataset quality by either analyzing the proportion of retained samples or by averaging samples’ mimic scores as an overall metric.
+We summarize the steps of Grad-Mimic in Algorithm 1 and detailed in Appendix A.
 5 Experiments
-We assess the effectiveness of mimic score within Grad-Mimic using two experimental setups across datasets of varying scales and domains.
+We evaluate the effectiveness of Grad-Mimic using two experimental setups across datasets of varying scales and domains.
 We first test in a controllable setting, adding noise into sample labels (Sec.
@@
 Our goals are to validate the following claims in both setups:
-  • Enhanced Model Performance: The reference model acts as a reliable guide. Training a new model with samples prioritized by mimic scores enhances performance.
-  • Accurate Sample Identification: Grad-Mimic effectively detects noisy samples, automates data selection, and complements existing filtering methods.
-  • Dataset Quality Assessment: Mimic scores and their associated filters offer reliable estimates of dataset quality, highly correlated with noise levels and performance gains.
-5.1 Simulating Mislabeled Samples
-Setups.
-We begin with a controlled experiment by adding various levels of label noise to six image classification datasets.
-They are DTD [47], Flowers102 [48], STL10 [49], OxfordIIIT Pet [50], CIFAR10 [51], and CIFAR100 [51].
-We fine-tune ViT-B/16 models [52] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples.
-We normalize mimic scores with a temperature of 0.5 and use a batch size of 32.
-We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last layer weights as the reference to navigate training on the noisy datasets.
-We detail more training configurations in Appendix C.1.
-After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and clustering methods using $k$-means and GMM, then aggregate filter outputs across training steps using Snorkel framework [31].
-Expected Results.
-We aim to validate that a pretrained reference model can serve as a strong guide for model training.
-Moreover, we expect that mimic scores obtained from Grad-Mimic can help identify low-value, mislabeled samples.
+  • Enhanced Data Efficiency: A pretrained reference model serves as a reliable guide, enabling mimic scores to prioritize samples that improve model performance.
+  • Accurate Sample Identification: Grad-Mimic can detect noisy samples, construct effective filters, and automate data selection.
+  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality.
 Full Fine-tuning DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
@@
 Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
-SGD 46.91 42.02 33.24 65.54 55.90 39.84 75.35 58.57 49.98 70.43 64.62 55.63 90.31 86.69 82.01 69.42 63.77 59.28 69.66 61.93 53.33
+Mini-batch SGD 46.91 42.02 33.24 65.54 55.90 39.84 75.35 58.57 49.98 70.43 64.62 55.63 90.31 86.69 82.01 69.42 63.77 59.28 69.66 61.93 53.33
 GraNd 29.73 8.83 15.43 10.80 2.78 27.92 78.39 52.32 64.30 56.83 45.38 45.43 87.56 83.79 80.04 18.65 8.83 7.84 46.99 33.66 40.16
@@
 Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
-SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
+Mini-batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
 GraNd 36.22 30.59 25.32 18.23 13.90 10.49 68.70 59.55 49.68 69.47 60.02 48.13 88.02 85.61 81.48 71.60 70.51 69.27 58.71 53.36 47.40
@@
 Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
-Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, leading to improved denoising and performance gains.
+Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and closing the performance gap to reference model.
 DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100
@@
 Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric.
-Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and identification methods.
+Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and proposed identification methods.
 Notably, its performance remains high and robust regardless of the noise level.
+5.1 Simulating Mislabeled Samples
+Setups.
+We begin with a controlled experiment by adding various levels of label noise to six image classification datasets.
+They are DTD [43], Flowers102 [44], STL10 [45], OxfordIIIT Pet [46], CIFAR10, and CIFAR100 [47].
+We fine-tune the ViT-B/16 model [48] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples.
+We normalize mimic scores with a temperature of 0.5 and use a batch size of 32.
+We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets.
+We detail more training configurations in Appendix B.1.
+After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster.
+Then we aggregate filter outputs across training steps using Snorkel framework [30].
 Figure 3: Mimic Score Distribution: Extracted from CIFAR100 dataset with a noise level of 0.5. The distributions clearly separate, aligning with the correctness of the labels.
+Expected Results.
+We aim to validate that a pretrained reference model’s weights can serve as a reliable guide. Moreover, we expect that mimic scores can help identify low-value, mislabeled samples.
 Baselines.
-We compare Grad-Mimic’s training method with other approaches that use gradient information for sample prioritization.
+We compare Grad-Mimic’s training method (Stage 1) with other approaches that use gradient information for sample prioritization.
 We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch.
 (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes.
-(iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of a random batch, then excludes outlier samples.
+(iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of another random batch, then excludes outlier samples.
 (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch.
+Furthermore, we compare Grad-Mimic with Rho-Loss [15], a reference model-based approach that prioritizes training on worth learning samples, as well as influence function-based methods.
+Results are presented in Appendix C.
+Grad-Mimic outperforms both techniques while additionally offering efficiency advantages.
 Stage 1 Results.
 We present Grad-Mimic’s two-stage results separately.
-The Stage 1 results, shown in Table 1, report testing accuracy across different methods and noisy datasets using full fine-tuning and linear probing.
-We evaluate Grad-Mimic’s robustness by averaging accuracy across datasets under different levels of label noise.
-Results show that Grad-Mimic effectively identifies and down-weights mislabeled samples, improving denoising and closing the performance gap to the reference model.
-Notably, in the full fine-tuning setup where we only target the last layer to mimic, we successfully guide the entire model weight updates, ultimately outperforming other methods.
-This also highlights the computational feasibility of our approach.
-These findings validate the idea that using the direction toward a better-performing model can guide model training and lead to enhanced performance.
+Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations.
+These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising and narrowing the performance gap with the reference model.
+Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other methods.
+This highlights its computational efficiency, as it avoids requiring sample gradients from the full set of model weights.
+Additionally, we explore the impact of reference model quality in Appendix C, further strengthening our approach.
+These findings validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
+Stage 2 Results.
+Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples.
+We use F1-scores to report the detection results in Table 2.
+Grad-Mimic accurately identifies mislabeled data across datasets and remains robust when facing different noise levels.
+We use CIFAR100 dataset as an example and visualize computed mimic score distribution at each training step in Figure 3.
+In this dataset, half of the labels are flipped.
+We can see their distributions clearly separate into two groups, confirming that mimic scores are informative signals to identify samples.
+Despite variations over training, these dynamics can be captured in our aggregation step.
+We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data.
+The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903.
+These demonstrate mimic score’s effectiveness for dataset quality estimation.
+5.2 Data Selection in Large-scale Web Datasets
+Setups.
+Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets.
+We use the small-scale and medium-scale DataComp datasets [17], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken.
+This means that our results might not be comparable to previous approaches on the DataComp Leaderboard.
+See here for details., respectively, to train CLIP model from scratch [2].
+In these datasets, noise is naturally presented in the data.
+We follow the training setup from DataComp [49, 17] and use publicly available pretrained CLIP model’s weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K.
+This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us.
+It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible.
+We use the final MLP layer in the text and image encoders respectively as our target to mimic.
+We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks.
+Other details can be found in Appendix B.2.
+In Stage 2, we use top-$k$ percent identification to create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
 Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
 ImageNet
-dist. shifts
+dis. shifts
 VTAB Retrieval
@@
 38 datasets ($\uparrow$)
-Standard Training 0.026 0.035 0.139 0.114 0.131
-0.03 0.026 0.035 0.153 0.115 0.136
+Vanilla Training 0.026 0.035 0.139 0.114 0.131
 0.05 0.026 0.035 0.147 0.114 0.135
-0.07 0.026 0.035 0.151 0.116 0.135
-0.3 0.026 0.034 0.154 0.112 0.137
 Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
-0.03 0.025 0.035 0.149 0.114 0.133
 0.05 0.026 0.037 0.162 0.118 0.146
-0.07 0.027 0.036 0.166 0.118 0.145
-0.3 0.026 0.035 0.161 0.114 0.144
 Small Grad-Mimic Last MLP Layer in Image Encoder 0.5 0.027 0.035 0.160 0.117 0.145
-Standard Training 0.171 0.148 0.253 0.217 0.254
+Vanilla Training 0.171 0.148 0.253 0.217 0.254
 Medium Grad-Mimic
@@
 0.05 0.169 0.151 0.262 0.216 0.258
-Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance.
-((a))
-((b))
-Figure 4: Mimic scores offer precise dataset quality estimation: Left figure (a) computes the portion of retained samples to evaluate dataset quality in simulation experiments.
-Right figure (b) takes the average on mimic scores in each filtered dataset for estimation.
-[S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
-Stage 2 Results.
-Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples.
-We use F1-scores to report the detection results in Table 2.
-Grad-Mimic accurately identifies mislabeled samples across all datasets and demonstrates robust performance when facing different noise levels.
-We use CIFAR100 dataset as an example and visualize mimic score distribution at each training step in Figure 3.
-In this dataset, half of the labels are flipped.
-We see their distributions clearly separate into two groups, confirming that mimic scores can serve as informative signals to identify samples.
-Despite variations over training, these dynamics can be effectively captured in our aggregation step.
-We use an ensemble filter from each dataset created by GMM clustering to estimate dataset quality using the proportion of retained data.
-The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903.
-This demonstrates mimic score’s effectiveness for dataset quality estimation.
-5.2 Data Selection in Large-scale Web Datasets
-Setups.
-We evaluate Grad-Mimic in a more challenging setting using million-scale web-crawled datasets.
-We use the small- and medium-scale DataComp datasets [16], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken.
-This means that our results might not be comparable to previous approaches on the DataComp Leaderboard.
-See here for details., respectively, for CLIP model pretraining [2].
-We follow the training setup from DataComp [53, 16] and use publicly available pretrained CLIP weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K.
-This pretrained model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us.
-It serves as a proxy to the ideal reference point for scenarios where training such powerful models is infeasible.
-We target to mimic the final MLP layer in the text and image encoders respectively.
-We evaluate model performance using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks.
-Additional training details can be found in Appendix C.2.
-In Stage 2, we use top-$k$ percent selection method to create the final filter, subsequently training a CLIP model on the curated dataset to validate filter’s effectiveness.
+Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance.
+Full table with various temperature settings is placed in Table 6 in Appendix  C.
 Expected Results.
-We anticipate mimic scores help large-scale training focus on high-value samples, while the derived scores will be instrumental in automating effective data curation.
+We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
 Baselines.
@@
 Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2].
-We see CLIP score as a strong baseline, which is prebuilt through heavy training using hundreds of millions of samples.
+We see CLIP score as a strong baseline, as it is prebuilt through heavy training using hundreds of millions of samples.
 Stage 1 Results.
 Table 3 presents pretraining results for CLIP models on both dataset scales.
-Grad-Mimic consistently outperforms standard training across all the temperature settings.
+Grad-Mimic consistently outperforms standard training across all temperature settings.
+Full results using different temperatures are provided in Appendix C.
 Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder.
-These results highlight Grad-Mimic’s success in using pretrained model weights as a guide to prioritize high-value samples to learn.
+These results highlight Grad-Mimic’s success in using online pretrained weights as a reference to prioritize high-value samples for learning.
 Scale Filtering Strategy Dataset Size ImageNet
@@
 CLIP Score (B/32 Top 30%) 3M 0.051 0.054 0.183 0.118 0.165
-CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M 0.047 0.053 0.174 0.106 0.163
-Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M 0.047 0.052 0.173 0.113 0.164
+CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M (0.3M $\downarrow$) 0.047 0.053 0.174 0.106 0.163
+Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M (0.4M $\downarrow$) 0.047 0.052 0.173 0.113 0.164
 No Filtering 101.9M 0.026 0.035 0.139 0.114 0.254
 Basic Filtering 30M 0.209 0.176 0.264 0.232 0.267
-Mimic Score Top 30% 29M 0.192 0.163 0.269 0.218 0.268
+Mimic Score Top 30% 29M (1M $\downarrow$) 0.192 0.163 0.269 0.218 0.268
 CLIP Score (B/32 Top 30%) 30M 0.275 0.229 0.332 0.246 0.321
-CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M 0.270 0.227 0.339 0.246 0.323
-CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M 0.274 0.226 0.334 0.243 0.323
-Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M 0.264 0.223 0.330 0.244 0.322
+CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M (1.9M $\downarrow$) 0.270 0.227 0.339 0.246 0.323
+CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M (2.3M $\downarrow$) 0.274 0.226 0.334 0.243 0.323
+Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M (4.7M $\downarrow$) 0.264 0.223 0.330 0.244 0.322
 Table 4: Stage 2 Results in DataComp Experiment: Mimic score-based filters perform better than basic filtering and complement CLIP score-based filters by removing low-value samples.
 Symbol $\setminus$” denotes the exclusion of curated datasets.
+((a))
+((b))
+Figure 4: Mimic scores offer precise dataset quality estimation: Left figure computes the portion of retained samples to evaluate dataset quality in simulation experiments.
+Right figure takes the average on mimic scores in each filtered dataset for estimation.
+[S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
 Stage 2 Results.
@@
 Results are presented in Table 4.
-Grad-Mimic successfully automates the selection process, outperforming the basic filtering strategy at both dataset scales.
-Additionally, we remove less useful samples identified by Grad-Mimic to improve CLIP score-based filter (B/32 Top 30%).
+Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales.
+Additionally, we remove low-value samples identified by mimic score to improve CLIP score-based filter (B/32 Top 30%).
 This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
 Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters.
-Besides, we include the top-ranked approach during our time of experimentation [54], which uses an orthogonal approach based on object detection models.
-By calculating the average mimic score in each curated dataset, we compare with their corresponding performance gains.
-The results, shown in Figure 4, reveal a strong alignment between performance gains and average mimic scores, with a Pearson correlation of 0.958, demonstrating that mimic scores can serve as a reliable metric for estimating dataset quality.
-Figure 5: The filtering logic behind mimic score aligns with handcrafted heuristics, particularly in caption length.
-Low-value samples (in the left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
+We include the top-ranked approach during our time of experimentation [50], which uses an orthogonal approach based on object detection models.
+We average samples’ mimic scores in each curated dataset and compare with their corresponding performance gains.
+The results, shown in Figure 4, reveal a positive alignment with a Pearson correlation of 0.958, demonstrating mimic score as a reliable metric for dataset quality estimation.
+Sample Analysis.
+We analyze high-value and low-value samples ranked by mimic score.
+In Figure 5, we compare token usage in captions.
+We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples.
+Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or incoherent images in the medium-scale dataset (see examples in Figure 1).
+These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
+We visualize the top and bottom 30% of samples along with their CLIP scores in Figure 6.
+High-value samples align with higher CLIP scores, while low-value ones are more concentrated in the lower range (half below 0.2).
+Furthermore, we calculate their correlation on random samples.
+We normalize both scores into the range 0 to 1 for clarity The observed positive correlation supports mimic scores in identifying high-quality samples.
+Figure 5: The filtering logic behind Grad-Mimic aligns with handcrafted heuristics, particularly in caption length.
+Low-value samples (left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
 ((a))
 ((b))
-Figure 6: Mimic score distribution and its relationship with CLIP score.
-Sample Analysis.
-We analyze high- and low-value samples ranked by mimic score.
-In Figure 5, we compare the token usage distribution in captions.
-We find that bottom-ranked samples often carry captions that are either too short (one or two words) or excessively long (exceeding the maximum token limit in CLIP’s tokenizer), unlike top-ranked samples.
-Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or visually incoherent images in the medium-scale dataset (see examples shown in Figure 1).
-These filtering patterns align with human intuition but are automatically captured by Grad-Mimic framework.
-We select the top 30% and bottom 30% of samples ranked by their mimic scores and visualize their distribution along with corresponding CLIP scores.
-In Figure 6, we observe high-value samples distributed towards higher CLIP scores, while low-value samples are more concentrated in lower CLIP score ranges (half of the samples are scored less than 0.2).
-Furthermore, we calculate their correlation using randomly selected samples.
-For a better visualization, we normalize both scores into the range 0 to 1.
-We see a positive correlation between them, supporting the reliability of mimic scores in identifying high-quality samples.
+Figure 6: Mimic score distribution and correlation to CLIP score.
 Filtering Strategy Random Selection Mimic Score Selection
@@
 We first apply various pre-built filters to curate datasets and train CLIP models on each.
-Then, we use their final layer weights as the reference and apply Grad-Mimic to train on the entire data pool.
-Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic scores appear in their curated dataset.
+Then, we target their final-layer weights to mimic and apply Grad-Mimic to train on the entire data pool.
+Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic score appear in their training dataset.
 We evaluate our approach against random selection using Jaccard similarity and percentage of overlap.
 As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling.
-We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking the final layer weights without direct access to their filtering steps.
+We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking final-layer weights without direct access to their filtering steps.
 6 Conclusion
-We introduce a new data quality metric called Mimic Score, designed to assess samples’ utility in the training process.
-We quantify samples by measuring the alignment of per-sample gradient and the vector toward a pretrained reference model.
-Building on Mimic Score, we develop Grad-Mimic to automate data selection process.
-Empirically, we demonstrate that Grad-Mimic effectively prioritizes identified samples to learn, resulting in enhanced model performance.
-Moreover, Mimic Score can serve as a reliable metric for creating effective filters and estimating dataset quality.
-Developed filters outperform human-designed ones and exhibit a strong correlation with performance gains.
+We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model training.
+We quantify samples by measuring the alignment between gradients and the vector toward pretrained reference weights.
+Using this, we develop Grad-Mimic to automate data selection.
+Empirically, Grad-Mimic effectively prioritizes samples for learning, improving data efficiency and model performance.
+Moreover, Mimic Score aids in creating effective filters and estimating dataset quality.
+Resulting filters outperform human-designed ones and exhibit a strong correlation with dataset quality and performance gains.
+7 Impact Statements
+Efficient data selection is increasingly important for training foundation models.
+We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and resulting model performance.
+We do not foresee explicit negative social impacts arising from Grad-Mimic.
+However, if the reference model used to evaluate sample utility is suboptimal, there remains a risk of refining low-quality datasets (we have an analysis of the impact of reference model quality in Appendix C).
+We believe that careful selection and validation of this reference model, a key area of future research, will minimize this risk and enable the creation of improved datasets.
 8 Acknowledgments
-We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, and John Cooper for their helpful feedback and valuable discussion.
-Appendix A Discussion
-A.1 Broader Impacts and Limitations.
-The importance of data selection has grown in building foundation models.
-We demonstrate that Grad-Mimic can effectively identify and exclude noisy or unhelpful samples.
-This leads to reduced computational and resource costs.
-Besides, Grad-Mimic provides the second benefit by enabling models to focus their training on high-value samples, which contributes to performance improvements.
-We do not foresee explicit negative impacts arising from Grad-Mimic.
-However, if the reference model used for sample selection is unreliable, there remains a risk of refining suboptimal datasets.
-To address this, incorporating human expert-designed filtering rules may be necessary to ensure the quality of curated datasets.
-Appendix B Grad-Mimic Algorithm
+We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, John Cooper, Harit Vishwakarma, Dyah Adila, Changho Shin, Satya Sai Srinath Namburi, Zhiqi Gao, Avi Trost, and Gabe Orlanski for their helpful feedback and valuable discussion.
+Appendix A Grad-Mimic Algorithm
 We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights.
-Our algorithm supports various training settings, such as supervised learning and self-supervised learning.
-We have validated both training scenarios through simulation experiments (Sec.
-5.1) and by training on web-crawled datasets with CLIP models (Sec.5.2).
-Moreover, these experiments evaluate Grad-Mimic performance across three different training configurations: full-parameter fine-tuning, linear probing, and training models from scratch.
+Our algorithm supports various training paradigms, including both supervised learning and self-supervised learning.
+We validate its effectiveness through controlled simulation experiments (Sec.
+5.1) and large-scale training of CLIP models on web-crawled datasets (Sec.
+5.2).
+Furthermore, our experiments encompass a range of training configurations, such as full-parameter fine-tuning, linear probing, and training models from scratch.
 Algorithm 1 Grad-Mimic (Stage 1)
 reference model $\theta_{\text{ref}}$, new model $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature $\tau$
-samples’ mimic scores $\overline{m}_{i,t}$, trained model $\theta_{T}$
+normalized mimic scores $\overline{m}_{i,t}$, a well-trained model $\theta_{T}$
 Initialize new weights $\theta_{0}$
@@
      uniformly sample a batch $\{s_{i}\}_{i=1}^{b}$ from $D$
-     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as guide
+     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as a guide
      for $i\in[b]$ do
-         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradients
-         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|_{2}}$triangleright compute mimic scores
+         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradient
+         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright compute mimic score
      end for
-     $\overline{m}_{:,t}=\frac{e^{-m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{-m_{j,t}/\tau}}$triangleright normalize mimic scores
+     $\overline{m}_{:,t}=\frac{e^{m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}$triangleright normalize mimic scores
      $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright reweight gradients
 end for
-Appendix C Experimental Details
-We provide more details about our training setups, and computational resources.
-C.1 Simulation Experiments
+Appendix B Experimental Details
+We provide more details about our training setups and computational resources.
+B.1 Simulation Experiments
 In the experiments of simulating mislabeled samples (Sec.
@@
 Results for other temperature values are presented in Table 7.
-To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets using identical configurations but with different random seeds for weight initialization.
+To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets but with different random seeds for weight initialization.
 These experiments were performed on a Nvidia Tesla A100.
+B.2 DataComp Experiments
+We adopt the training setup from DataComp [49, 17] and train CLIP models from scratch using contrastive objective over image-caption pairs.
+Each CLIP model is trained for 5 epochs with a batch size of 4096.
+The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset.
+These experiments run on 8 Nvidia Tesla A100s.
+We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders.
+Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [17] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5).
+We report results at $\tau=0.5$ and $\tau=0.05$ in the main body in Table 3, but full results are presented in Table 6.
+Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
+ImageNet
+dist. shifts
+VTAB Retrieval
+Average over
+38 datasets ($\uparrow$)
+Vanilla Training 0.026 0.035 0.139 0.114 0.131
+0.03 0.026 0.035 0.153 0.115 0.136
+0.05 0.026 0.035 0.147 0.114 0.135
+0.07 0.026 0.035 0.151 0.116 0.135
+0.3 0.026 0.034 0.154 0.112 0.137
+Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
+0.03 0.025 0.035 0.149 0.114 0.133
+0.05 0.026 0.037 0.162 0.118 0.146
+0.07 0.027 0.036 0.166 0.118 0.145
+0.3 0.026 0.035 0.161 0.114 0.144
+Small Grad-Mimic Last MLP Layer in Image Encoder 0.5 0.027 0.035 0.160 0.117 0.145
+Vanilla Training 0.171 0.148 0.253 0.217 0.254
+Medium Grad-Mimic
+Last MLP Layer
+in Image Encoder
+0.05 0.169 0.151 0.262 0.216 0.258
+Table 6: Full Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance across all temperature settings.
+Appendix C Ablation Studies
 Temperature $\tau$ DTD Flower102 Oxford-IIIT Pet CIFAR10 CIFAR100
-SGD 52.8 35.6 87.9 93.2 76.2
+Mini-Batch SGD 52.8 35.6 87.9 93.2 76.2
 Grad-Match 53.1 32.1 87.4 93.0 76.1
@@
 Grad-Mimic 0.6 56.0 46.4 89.3 94.2 77.3
-Table 6: Stage 1 Results in Simulation Experiment with Different Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
-C.2 DataComp Experiments
-We adopt the training setup from DataComp [53, 16].
-CLIP models are trained from scratch using a contrastive objective over image-caption pairs.
-Each model is trained for 5 epochs with a batch size of 4096.
-The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset.
-These experiments ran on 8 Nvidia Tesla A100s.
-We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders.
-The layers used are “visual.transformer.resblocks.11.mlp.c_fc.weight” and “text.transformer.resblocks.11.mlp.c_fc.weight”.
-Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [16] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5).
-Results are presented in Table 3.
-Appendix D Ablation Studies
+Table 7: Stage 1 Results in Simulation Experiment using Various Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
 The Choice of Temperature.
-We present Grad-Mimic results with different temperature values and compare them to baseline methods (SGD and Grad-Match [24]).
+We test Grad-Mimic using different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [24]).
 The results in the simulation experiments are presented in Table 7.
-We set the noise level to 0.3 and fine-tune ViT model under linear probing configuration.
-Grad-Mimic outperforms the baseline methods across all temperature settings.
+In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration.
+Grad-Mimic outperforms the baseline methods in all temperature settings.
 Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
+DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
+Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
+Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
+Mini-Batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
+Rho-Loss 54.10 52.39 48.46 41.55 35.71 30.02 97.10 96.83 96.49 88.53 87.79 86.86 94.07 93.83 93.82 76.82 76.05 75.93 75.36 73.77 71.93
+Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
+Table 8: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing memory usage and computational overhead.
+Comparison to Reference Model-Based Approaches.
+We compare Grad-Mimic to Rho-Loss [15], a method that prioritizes samples for learning based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on the pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$.
+We follow our simulation experiment setup and evaluate the performance of fine-tuned models on the testing datasets.
+As shown in Table 8, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels.
+Furthermore, Grad-Mimic offers efficiency advantages.
+Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes, Grad-Mimic computes mimic scores directly from the reference model’s weights.
+In our experiments, we demonstrate that even using only a subset of the weights (e.g., the last-layer weights) is sufficient, further reducing memory usage and improving computational efficiency.
+Figure 7: Grad-Mimic can outperform influence function-based methods: Using model weights as a guide offer advantages on their effectiveness and greater accessibility.
+Comparison to Influence Function-based Approaches.
+Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [52, 53], but their computational cost often limits practical use.
+Several studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a validation dataset [54, 55].
+We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from validation dataset.
+We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from each class (1, 5, 10, 15, and 20).
+Results in Figure 7 show that while larger validation sets improve data efficiency by enabling model to focus on the influential samples, Grad-Mimic—using pretrained reference weights and their induced direction—consistently outperforms validation dataset-based influence approximations.
+Moreover, relying on a labeled validation set risk introducing an unwanted dependency.
+For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
+Figure 8: Grad-Mimic can be robust to the quality of reference model.
+Impact of Reference Model Quality.
+Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available.
+To simulate such scenarios, we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights.
+By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
+We present this analysis in Figure 8.
+As expected, increasing noise variance degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic.
+However, we observe that even under significant noise, models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself.
+This highlights that while a high-quality reference model is ideal, even a weaker proxy can provide valuable guidance and improve training outcomes over traditional baselines.

### v3

@@
-Evaluating Sample Utility for Data Selection by Mimicking Model Weights
+Evaluating Sample Utility for Efficient Data Selection by Mimicking Model Weights
 Abstract
-Foundation models are trained on large-scale web-crawled datasets, which often contain noise, biases, and irrelevant information.
-This motivates the use of data selection techniques, which can be divided into model-free variants—relying on heuristic rules and downstream datasets—and model-based, e.g., using influence functions.
-The former can be expensive to design and risk introducing unwanted dependencies, while the latter are often computationally prohibitive.
-Instead, we propose an efficient, model-based approach using the Mimic Score, a new data quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model.
-It relies on the alignment between gradients and a target direction induced by the reference model.
-Using the derived Mimic Scores, we develop Grad-Mimic, a framework that prioritizes samples for learning, creates effective filters, and automates data selection.
-Empirically, using Mimic Scores to guide training improves data efficiency, results in consistent performance gains across six image datasets, and includes enhancements to CLIP models.
-Moreover, Mimic Score-based filters improve upon existing filtering methods, e.g., cutting 4.7 million samples to train better CLIP models while offering accurate estimation of training dataset quality.
+Multimodal models are trained on large-scale web-crawled datasets, which often contain noise, bias, and irrelevant information.
+This motivates the use of data selection techniques, which can be divided into model-free variants, relying on heuristic rules and downstream datasets, and model-based approaches, such as those using influence functions.
+The former can be expensive to design and risks introducing unwanted dataset dependencies, while the latter are often computationally prohibitive.
+In this work, we propose an efficient, model-based approach using the Mimic Score, a new data-quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model.
+Our method relies on measuring alignments between training gradients and a target direction induced by this reference model.
+Building on the derived mimic scores, we develop Grad-Mimic: a framework that prioritizes samples to learn, estimates overall sample utility, and creates effective filters.
+Empirically, using mimic scores to guide training improves data efficiency, accelerates convergence, yields consistent performance gains across six image datasets, and enhances CLIP models with 20.7% fewer training steps.
+Moreover, mimic score-based filters complement existing filtering methods, e.g., training improved CLIP models with 4.7 million fewer samples while offering accurate estimation of dataset quality.
 1 Introduction
-Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], and GPT-4 [3].
-These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content.
+Large-scale web-crawled datasets are fundamental to the success of modern multimodal models such as AIMv2 [1], ALIGN [2], OpenAI CLIP [3], and SigLIP [4].
+These datasets provide vast quantities of information—but inevitably contain noise, biases, and irrelevant content.
 To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [5, 6].
-For example, the FineWeb dataset [6], with 15 trillion tokens, is refined through eight carefully designed filtering steps to ensure high model performance.
-Existing data selection methods have several limitations.
-We can divide these into two categories: those that are independent of a model being trained (or using pretrained models) and those that rely on such a model.
-Model-free approaches, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [7, 8] often require expensive experimentation to establish selection rules and add unwanted dependencies on additional datasets.
-Furthermore, these methods can fail to provide fine-grained insights into individual samples and overlook sample utility in training, resulting in suboptimal performance.
-In contrast, model-based techniques, e.g., training specialized filtering networks  [9, 10, 11, 12] or using data influence functions  [13, 14], introduce complexity and are often computationally expensive.
+For example, the FineWeb dataset [7], containing 15 trillion text tokens, undergoes eight carefully designed filtering steps to ensure high model performance.
+Current selection strategies; however, are far from satisfactory.
+We can broadly divide them into two categories: model-free methods, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [8, 9] often require costly trial-and-error to establish selection rules and add unwanted dependencies on other datasets.
+These methods also lack the granularity needed to assess individual utility, leading to suboptimal training outcomes.
+In contrast, model-based techniques, such as those building specialized filtering networks [10, 11, 12, 13], scoring samples through a pretrained reference model’s loss [14, 15], or using influence functions  [16, 17, 18, 19, 20], increase the complexity of the pipeline and are often computationally expensive.
+These shortcomings motivate the need for a simpler, dataset-agnostic, compute-efficient way to curate high-quality datasets.
 Figure 1: High-value vs.
-low-value samples identified by the Mimic score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their Mimic scores.
-Each image is accompanied by its caption and corresponding CLIP score (higher scores suggest better quality).
+low-value samples identified by the Mimic score: Randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their mimic scores.
+Each image includes its caption and CLIP score (higher indicates better quality).
 Mimic scores align closely with CLIP scores.
-High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
-To address these limitations, we introduce Mimic Score, a new data quality metric that assesses sample’s contribution to weight updates using reference model’s weights.
-This is an efficient model-based technique.
-We show that by leveraging pretrained model weights (located in a more optimal part of the weight space), we bypass the need for downstream datasets or specialized model training and can use it as our selection guide.
-Our approach is based on the alignment between each sample’s negative gradient and a direction pointing toward the reference model in the weight space.
-We argue that samples that potentially pull the model in undesirable directions—thereby misguiding weight updates—should be considered low-value and ruled out.
-Leveraging model weights to evaluate sample utility offers several advantages.
-First, with the growing availability of well-trained models, adopting model weights has become more accessible than ever.
-Second, unlike other reference model-based approaches that require loading a second model and performing full inference passes [15, 16], our method scores each sample directly from the weights, improving efficiency.
-Building on the mimic score, we develop Grad-Mimic, a data selection framework that operates in two stages:
-  • Stage 1: During training, Grad-Mimic prioritizes which samples to learn based on their mimic scores, improving data efficiency while steering the model toward the reference model’s weight space.
-  • Stage 2: After training, Grad-Mimic uses mimic scores to evaluate each sample’s utility across training steps.
-These assessments are then aggregated to create an ensemble filter, which can further refine the dataset.
-We validate Grad-Mimic through diverse experiments.
-First, in a controlled setting with varying levels of label noise, Grad-Mimic accurately identifies mislabeled samples and estimates training dataset quality.
-During training, it effectively down-weights undesirable samples and improves data efficiency across six image datasets.
-Next, we test it on large-scale web-crawled multimodal data by training CLIP models from scratch [2] on 10M and 100M samples from DataComp  [17].
-Using publicly available pretrained weights as our reference, mimic scores effectively navigate training.
-The resulting filter performs better than human-designed ones and can complement CLIP score-based filters [18], lifting CLIP model performance with less training data.
+High-value samples generally have detailed captions and coherent visuals, while low-value ones carry short captions and misaligned content.
+We introduce the Mimic Score, a new data-quality metric that quantifies each training sample’s contribution to effective weight updates.
+We argue that samples whose gradients potentially pull the model in undesirable directions—and thus misguide weight updates—should be down-weighted or discarded.
+To establish a desirable weight update, we leverage the direction pointing toward the weights of a pretrained reference model111 Note that our goal is to select high-quality data, not to reproduce the reference model.
+While insight into reference model’s training data would be valuable, we assume it is inaccessible—a common situation with many modern open-source models.
+, located in a more optimal part of the weight space, as our proxy.
+The mimic score is computed for each sample by measuring the alignment between its negative gradient and the direction induced by this reference model.
+A high alignment score indicates a sample can steer the model toward this preferred region, while a low one would divert the model.
+Using model weights to evaluate sample utility offers multiple practical benefits.
+First, the growing supply of well-trained models makes this approach more accessible than ever.
+Second, unlike methods that derive selection guides using the gradients on noise-free validation datasets [18, 19, 20], mimic score avoids extra validation computations, bypassing the difficulties of curating such dataset and ensuring its quality.
+Finally, many reference model-based techniques compute excess loss through additional full inference [14, 15].
+Our mechanism instead is a much simpler operation based on the discrepancy of model weights, which can substantially reduce compute overheads.
+Building on the mimic score, we develop Grad-Mimic, a two-stage data-selection framework:
+  • Online Batch Re-weighting. During training, Grad-Mimic computes mimic scores on the fly and prioritizes samples to learn, enhancing data efficiency and accelerating convergence.
+  • Offline Sample Selection.
+After training, Grad-Mimic aggregates sample’s mimic scores across training steps to estimate an overall sample utility.
+Such combined estimate serves as a reliable filter, enabling the construction of a smaller, higher-quality dataset for future use.
+We validate the capabilities of Grad-mimic in a broad range of scenarios.
+For mislabeled sample detection, we show in a controlled setting, Grad-Mimic accurately identifies noisy labels and estimates overall dataset quality.
+By down-weighting undesirable samples during training, it enhances data efficiency and model performance across six image datasets.
+For large-scale data curation, we scale our testbed to web-crawled multimodal data by training CLIP models from scratch [3] on 10M and 100M samples [21].
+Using publicly released pretrained weights as our reference, mimic scores effectively navigate training, reducing 20.7% training steps to convergence.
+The resulting filter outperforms human-designed ones and complements CLIP score-based filters [22], lifting CLIP model performance with 4.7M fewer training samples.
 We summarize our contributions as follows:
-  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between gradients and the vector toward the reference model in the weight space.
-  • Automated Data Selection Framework: Using the mimic score, we present Grad-Mimic, a framework that effectively identifizes useful samples to improve data efficiency and automates effective data selection.
-  • Dataset Quality Assessment: The mimic score, along with its resulting filters, improves existing filtering strategies and provides reliable estimation of dataset quality.
+  • New Data-quality Metric: We introduce Mimic Score to assess sample utility by measuring the alignment between training gradients and a vector toward the reference model in the weight space.
+  • Efficient Data Selection Framework: Built on Mimic Score, we present Grad-Mimic, a framework that efficiently prioritizes valuable samples to learn, enhances data efficiency, accelerates convergence, and automates effective data selection.
+  • Dataset Quality Assessment: Mimic Score, along with its ensemble filters, provides reliable estimates of dataset quality and model performance gains and improves existing filtering strategies.
 2 Related Work
-Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients.
-Then it adapts alignment scores to reweight gradients that contributed to the weight updates.
-In Stage 2, derived alignment scores—mimic scores—are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality estimation, and training dataset discovery.
+Our work explores using the weights of a reference model to evaluate sample utility, intersecting three key areas: (i) data selection, (ii) data curation for multimodal models, and (iii) weak supervision.
+Figure 2: Grad-Mimic Two-stage Workflow: During training, Grad-Mimic measures alignments using the vector toward the reference model’s weight space against each sample’s negative gradient.
+It then adapts alignment scores to re-weight gradients for weight updates.
+After training, the derived alignment scores, mimic scores, are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality assessment, and training dataset discovery.
 Data Selection.
 Data selection techniques can be categorized into group- and sample-level approaches.
-Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [19, 20, 21].
-Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and filter out unhelpful samples.
-Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training [24, 25].
-These methods can struggle with large amounts of noise in batch gradients.
-Grad-Mimic overcome this with the help of a reference model.
+Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [23, 24, 25, 26].
+Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and rule out undesired samples.
+Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [27], analyzing gradient similarities to other training batches [28], selecting key samples that can capture full training [29, 30], and using gradients computed on noisy-free validation sets [18, 19, 20].
+These methods; however, are sensitive to noisy batch gradients and often incur extra computation.
+Grad-Mimic sidesteps both issues with the help of a reference model.
 Data Curation for Multimodal Models.
-Multimodal models rely on large-scale web datasets [26, 17, 27], which are often naturally noisy and require careful curation.
+Multimodal models rely on large-scale web datasets [31, 21, 32]; however, quantity comes at a price: unconstrained crawling amplifies noise and requires careful curation.
 Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [8, 9], using semantic deduplication [33], developing specialized filtering networks [10, 11, 12, 13], and using data influence functions [16, 17].
-While effective, they require access to target datasets, introduce training complexities, and can be expensive in computation.
-Grad-Mimic overcomes these and offers a more efficient alternative: using public available model weights to identify useful samples.
+While effective, they require access to other datasets, introduce training complexities, or demand substantial compute.
+Grad-Mimic offers a more efficient alternative: using public available models and the discrepancy in model weights to identify useful samples.
 Weak Supervision.
-Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [29, 30, 31, 32].
+Weak supervision is a framework to construct labeled datasets by combining multiple noisy label estimates [34, 35, 36, 37].
 These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [38, 39], often encoded as labeling functions.
-The outputs of these labeling functions are modeled and then aggregated to produce final labels [29].
-Weak supervision has demonstrated success across diverse domains [35, 36, 37, 38, 39].
-Most prior works focus on weak label aggregation to construct datasets, Grad-Mimic leverages weak supervision for a new purpose: modeling sample utility assessments for a combined filtering decision.
+Labeling functions’ outputs are modeled and aggregated to produce probabilistic labeling decisions [34].
+Weak supervision has demonstrated success in diverse domains [40, 41, 42, 43, 44].
+Most prior works focus on weak label aggregation to construct datasets.
+Grad-Mimic leverages it for a new purpose: modeling each sample’s utility assessments for a combined filtering decision.
 3 Evaluating Sample Utility
-We quantify samples by their contributions to the learning process.
+Our goal is to quantify training sample contributions to learning process when we (i) have a model to be trained and (ii) access a pretrained reference model as proxy.
 Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value.
-We start with setup and notation, then explain how the scoring metric is derived and used in our data selection framework, Grad-Mimic.
+We start with setup and notation, then explain how an intermediate scoring metric is derived and used inside our data selection framework, Grad-Mimic (Sec.
+4).
 Notation.
@@
 Each sample $s_{i}$ can be expressed as $(x_{i},y_{i})$, where noise may be present either in the instance $x_{i}$ or in the label $y_{i}$.
-The empirical loss across samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$.
-A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
-3.1 Mimic Score Calculation
-To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our guide.
-These reference weights, denoted by $\theta_{\text{ref}}$, reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$.
+The empirical loss across all the samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$.
+A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\frac{1}{b}\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
+Mimic Score Calculation.
+To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our proxy guide.
+These reference weights, denoted by $\theta_{\text{ref}}$, are assumed reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$.
 We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$).
 These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [45, 46].
-Reference weights can be obtained in two ways: by training a model on the dataset to achieve a desired performance level, if resources permit, or more efficiently, by using publicly available well-trained models, which are now more accessible and can eliminate the need for direct data access or possible costly training.
+Reference weights can be obtained either by training a model on a (hold-out) dataset until it meets the desired performance [15, 14], if resource permits, or, more efficiently, by using an existing well-trained public model, which sidesteps both data-access constraints and costly training.
 The vector pointing toward reference model’s weights, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$.
@@
 […]
-This alignment score, named the mimic score, reflects how much a sample can drive the model closer to the reference model.
-A sample having a lower mimic score suggests it has limited utility in guiding model updates, making it a potential candidate for exclusion from future training.
-4 Grad-Mimic: Data Selection Framework
-Building on this simple but effective metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn and then identifies useful samples to design an effective filter.
-The complete Grad-Mimic workflow is illustrated in Fig.
+This alignment score, named mimic score, reflects how much a sample can drive the model closer to a preferred weight space.
+A sample having a lower mimic score suggests it has limited utility in guiding effective weight updates, making it a potential candidate for exclusion from future training.
+4 Data Selection Framework
+Building on this simple data-quality metric, we package it into Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn (Sec.
+4.1) and then combines sample utilities to design an effective ensemble data filter (Sec.
+4.2).
+A complete Grad-Mimic workflow is in Fig.
 2.
-4.1 Stage 1: Reweighted Gradient Updates
-We first use mimic scores to aid model training by re-weighting sample gradients.
-Unlike standard gradient descent, which assigns equal weight to all the samples in the mini-batch, Grad-Mimic uses mimic scores to amplify helpful samples and down-weight unhelpful ones.
+4.1 Stage 1: Online Batch Re-weighting.
+We first use mimic scores to aid model training through improving data efficiency by re-weighting.
+Unlike gradient descent, which assigns uniform weight to mini-batch samples, Grad-Mimic computes mimic scores on the fly to amplify helpful samples and down-weight unhelpful ones.
 To achieve this, each sample’s mimic score is first normalized using the softmax function with a temperature parameter, $\tau$.
@@
 The temperature $\tau$ controls the sensitivity of sample reweighting, allowing us to adjust how sharply the model prioritizes samples.
-A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher temperature encourages Grad-Mimic converges to standard gradient descent.
-4.2 Stage 2: Automated Filter Design
-The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, creating a combined filter based on sample utility assessments.
-4.2.1 Sample Identification
+A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher one encourages Grad-Mimic converges to gradient descent.
+Theoretical Analysis. We provide an analysis for effective batch reweighting in Appendix A. In particular, we study conditions where Grad-Mimic converges faster than standard gradient descent.
+4.2 Stage 2: Offline Sample Selection.
+The second stage of Grad-Mimic converts computed mimic scores, estimated sample utilities, into a unified assessment, enabling us to curate a higher-quality training dataset with developed data filter.
+Mimic Score Binarization.
 Grad-Mimic first gathers normalized mimic scores for each sample at every training step.
-These scores indicate the sample contributions over training, which allows us to decide whether to retain or discard a sample.
-Grad-Mimic supports several sample identification methods:
-  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
-  • 1D Clustering: Samples are categorized into two groups using clustering such as $k$-means clustering [42] or Gaussian Mixture Models (GMMs), allowing more automated identification based on their assigned cluster.
+These indicate their intermediate contributions to learning.
+We convert continuous scores into binary retain/discard decisions and explore three schemes:
+  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. A natural choice can be set as $1/b$ (indicating greater than uniform).
+  • 1D Clustering: Samples are categorized into two groups using clustering (e.g., $k$-means [47] or Gaussian mixture), allowing unsupervised identification based on their assigned clusters.
   • Top-k Percent: Samples are ranked by their mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
-After identification, normalized mimic scores are binarized, and samples are categorized into “retain” or “discard” at each training step.
-We treat these assigned groups as each training step’s votes and then combine for a final filter.
-4.2.2 Filter Output Aggregation
-Although using the assessment at a training step can refine the dataset, relying on such isolated filters risks overlooking training dynamics.
-Moreover, evaluating a sample’s utility across all training stages may yield complementary signals.
-To address this, Grad-Mimic uses weak supervision techniques [29, 30, 31, 32], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps.
-We start by learning a model to evaluate the reliability of each step’s assessment [31].
-Once established, this model aggregates filtering decisions into a high-quality ensemble filter.
-Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics.
-We employ the Snorkel framework [30], a widely used method in the weak supervision community.
-Ultimately, we can use this ensemble filter to select high-value samples for a refined dataset.
-Moreover, we can use it to estimate training dataset quality by either analyzing the proportion of retained samples or by averaging samples’ mimic scores as an overall metric.
-We summarize the steps of Grad-Mimic in Algorithm 1 and detailed in Appendix A.
+Decision Aggregation.
+A filtering decision made at a single step can be noisy and ignore training dynamics.
+To address this, we treat each step’s binarized assessments as weak votes and use weak supervision techniques [34, 35, 36, 37] to combine filtering decisions across training steps.
+We start by learning a generative model to estimate the reliability of each step’s assessments [36].
+Once established, this model produces probabilistic aggregated decisions.
+This consensus mitigates assessment noise and captures how a sample’s utility evolves.
+Empirically, we employ the aggregation step from Snorkel framework [35], a widely-used method in the weak supervision community.
+Ultimately, the ensemble filtering decisions yield a curated subset that can be reused in later training runs.
+Moreover, we can use their statistics to estimate training dataset quality by either analyzing retention rate (fraction of samples retained) or averaged mimic score.
+We discuss weak supervision setups and summarize algorithms in both online and offline stages in Appendix B.
 5 Experiments
-We evaluate the effectiveness of Grad-Mimic using two experimental setups across datasets of varying scales and domains.
-We first test in a controllable setting, adding noise into sample labels (Sec.
-5.1), followed by evaluations on large-scale web-crawled datasets (Sec.
+We validate Grad-Mimic’s effectiveness using two experimental setups across diverse scales and domains.
+First, we test in a controlled setting, adding noise sample labels (Sec.
+5.1), then evaluate on large-scale web-crawled datasets (Sec.
 5.2).
-Our goals are to validate the following claims in both setups:
-  • Enhanced Data Efficiency: A pretrained reference model serves as a reliable guide, enabling mimic scores to prioritize samples that improve model performance.
-  • Accurate Sample Identification: Grad-Mimic can detect noisy samples, construct effective filters, and automate data selection.
-  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality.
+Our goals are to confirm key claims in both settings:
+  • Enhanced Data Efficiency: The weights of pretrained reference model serve as reliable guide. Using derived mimic scores to prioritize samples improves data efficiency and speed up convergence.
+  • Accurate Sample Identification: Grad-Mimic precisely detects noisy samples, constructs effective ensemble filter, and automates data selection for high-quality datasets.
+  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality and predict model performance gains.
+5.1 Simulating Mislabeled Samples
+Setups.
+We begin with a controlled experiment by adding various levels of label noise to six image classification datasets.
+They are DTD [48], Flowers102 [49], STL10 [50], OxfordIIIT Pet [51], CIFAR10, and CIFAR100 [52].
+We fine-tune the ViT-B/16 model [53] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples.
+We normalize mimic scores with a temperature of 0.5 and use a batch size of 32.
+We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets.
+We detail more training configurations in Appendix C.
+Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and denoising capability.
 Full Fine-tuning DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
-Ref. Model 65.59 91.01 93.20 82.88 94.89 81.34 84.82
 Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
@@
 Linear Probing DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
-Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
 Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
@@
 Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
-Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and closing the performance gap to reference model.
+Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score.
+Grad-Mimic reliably detects mislabeled samples across datasets and binarization methods, maintaining strong performance despite varying noise levels.
 DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100
@@
 GMM 97.85 96.72 95.68 98.39 96.96 94.21 97.96 97.16 95.69 98.04 97.07 95.86 95.70 92.89 88.62 95.86 94.31 92.55
-Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric.
-Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and proposed identification methods.
-Notably, its performance remains high and robust regardless of the noise level.
-5.1 Simulating Mislabeled Samples
-Setups.
-We begin with a controlled experiment by adding various levels of label noise to six image classification datasets.
-They are DTD [43], Flowers102 [44], STL10 [45], OxfordIIIT Pet [46], CIFAR10, and CIFAR100 [47].
-We fine-tune the ViT-B/16 model [48] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples.
-We normalize mimic scores with a temperature of 0.5 and use a batch size of 32.
-We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets.
-We detail more training configurations in Appendix B.1.
-After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster.
+After training, we binarize samples’ mimic scores in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster.
 Then we aggregate filter outputs across training steps using Snorkel framework [35].
 Figure 3: Mimic Score Distribution: Extracted from CIFAR100 dataset with a noise level of 0.5. The distributions clearly separate, aligning with the correctness of the labels.
-Expected Results.
-We aim to validate that a pretrained reference model’s weights can serve as a reliable guide. Moreover, we expect that mimic scores can help identify low-value, mislabeled samples.
+Expected Results. We aim to validate that a pretrained reference model’s weights can serve as a reliable selection guide, and we expect mimic scores can help identify low-value, mislabeled samples.
 Baselines.
-We compare Grad-Mimic’s training method (Stage 1) with other approaches that use gradient information for sample prioritization.
+We compare Grad-Mimic’s Stage 1 online method with other approaches that use gradient information for sample prioritization.
 We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch.
@@
 (iv) Grad-Match [29] adapts gradients by solving a subset selection problem to identify key samples within each batch.
-Furthermore, we compare Grad-Mimic with Rho-Loss [15], a reference model-based approach that prioritizes training on worth learning samples, as well as influence function-based methods.
-Results are presented in Appendix C.
-Grad-Mimic outperforms both techniques while additionally offering efficiency advantages.
+Furthermore, we compare Grad-Mimic with Rho-Loss [14], a reference model-based approach that prioritizes training on worth learning samples, as well as approximated influence function-based methods.
+Results are presented in Appendix D.
+Grad-Mimic outperforms both techniques while additionally offering efficiency advantages (See Appendix E).
 Stage 1 Results.
@@
 Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations.
-These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising and narrowing the performance gap with the reference model.
-Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other methods.
-This highlights its computational efficiency, as it avoids requiring sample gradients from the full set of model weights.
-Additionally, we explore the impact of reference model quality in Appendix C, further strengthening our approach.
-These findings validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
+These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising capability.
+Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other baseline methods.
+Beyond denoising improvements, Grad-Mimic enjoys many efficiency benefits than others, including faster convergence, lower compute runtime, and reduced GPU memory usage.
+We place these analysis in Appendix E.
+Additionally, we explore the impact of reference model quality in Appendix F, further strengthening our approach.
+These findings all validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
 Stage 2 Results.
@@
 We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data.
-The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903.
+The results, displayed in Figure 4 (middle), highly correspond to the presence of label noise, with a Pearson correlation of -0.903.
 These demonstrate mimic score’s effectiveness for dataset quality estimation.
-5.2 Data Selection in Large-scale Web Datasets
-Setups.
-Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets.
-We use the small-scale and medium-scale DataComp datasets [17], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken.
-This means that our results might not be comparable to previous approaches on the DataComp Leaderboard.
-See here for details., respectively, to train CLIP model from scratch [2].
-In these datasets, noise is naturally presented in the data.
-We follow the training setup from DataComp [49, 17] and use publicly available pretrained CLIP model’s weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K.
-This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us.
-It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible.
-We use the final MLP layer in the text and image encoders respectively as our target to mimic.
-We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks.
-Other details can be found in Appendix B.2.
-In Stage 2, we use top-$k$ percent identification to create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
+Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance.
+Full table with various temperature settings is placed in Table 8 in Appendix  F.
 Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
@@
 0.05 0.169 0.151 0.262 0.216 0.258
-Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance.
-Full table with various temperature settings is placed in Table 6 in Appendix  C.
-Expected Results.
-We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
+5.2 Data Selection in Large-scale Web Datasets
+Setups.
+Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets.
+We use the small-scale and medium-scale DataComp datasets [21], which contain approximately 10 million and 100 million image-caption pairs [^2] We identified that some URLs provided by DataComp dataset are now broken.
+This means that our results might not be comparable to previous approaches on the DataComp Leaderboard.
+See here for details., respectively, to train CLIP model from scratch [3].
+In these datasets, noise is naturally presented in the data.
+We follow the training setup from DataComp [54, 21] and use publicly available pretrained CLIP model’s weights as our reference  [^3] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K.
+This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us.
+It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible.
+We use the final MLP layer in the text and image encoders respectively as our target to mimic.
+We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks.
+Other implementation details can be found in Appendix C.
+In Stage 2, we use top-$k$ percent approach to binarize scores and create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
+Expected Results. We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
 Baselines.
-In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally.
-After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [17]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [18]: selecting top-$k$ percent samples based on embedding alignment between images and captions.
+For the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally.
+For Stage 2 comparison, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [21]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [22]: selecting top-$k$ percent samples based on embedding alignments between images and captions.
 Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [3].
@@
 Table 3 presents pretraining results for CLIP models on both dataset scales.
-Grad-Mimic consistently outperforms standard training across all temperature settings.
-Full results using different temperatures are provided in Appendix C.
+Grad-Mimic consistently outperforms standard training at all temperature settings.
+Full results using various temperatures are provided in Appendix F.
 Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder.
-These results highlight Grad-Mimic’s success in using online pretrained weights as a reference to prioritize high-value samples for learning.
+Figure 4 (left) presents the learning curves for CLIP models trained with Grad-Mimic versus the vanilla method.
+By re-weighting samples based on mimic scores in weight updates, Grad-Mimic achieves faster convergence, requiring 20.7% fewer steps.
+These findings highlight Grad-Mimic’s success in leveraging online pretrained weights to prioritize high-value samples for efficient training.
+Table 4: Stage 2 Results in DataComp Experiment: Mimic score-based filters perform better than basic filtering and complement CLIP score-based filters by removing low-value samples.
+Symbol $\setminus$” denotes the exclusion of curated datasets.
 Scale Filtering Strategy Dataset Size ImageNet
@@
 Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M (4.7M $\downarrow$) 0.264 0.223 0.330 0.244 0.322
-Table 4: Stage 2 Results in DataComp Experiment: Mimic score-based filters perform better than basic filtering and complement CLIP score-based filters by removing low-value samples.
-Symbol $\setminus$” denotes the exclusion of curated datasets.
-((a))
-((b))
-Figure 4: Mimic scores offer precise dataset quality estimation: Left figure computes the portion of retained samples to evaluate dataset quality in simulation experiments.
-Right figure takes the average on mimic scores in each filtered dataset for estimation.
-[S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
+Figure 4: The left figure shows mimic score-guided training converges faster than vanilla baseline.
+The middle one uses retention rate to estimate training dataset quality in simulation experiments.
+The right figure takes the average on mimic scores in each curated dataset for performance gain estimation.
+[S/M] denotes the scale of DataComp dataset, small [S] or medium [M], that the filter is designed on.
 Stage 2 Results.
@@
 Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales.
-Additionally, we remove low-value samples identified by mimic score to improve CLIP score-based filter (B/32 Top 30%).
-This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
+Additionally, we remove low-value samples ranked by mimic score to improve CLIP score-based filter (B/32 Top 30%).
+This complementary approach not only enhances model performance but also improves data efficiency by reducing training set size, specifically removing 4.7 million samples in the medium-scale dataset.
 Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters.
 We include the top-ranked approach during our time of experimentation [55], which uses an orthogonal approach based on object detection models.
-We average samples’ mimic scores in each curated dataset and compare with their corresponding performance gains.
-The results, shown in Figure 4, reveal a positive alignment with a Pearson correlation of 0.958, demonstrating mimic score as a reliable metric for dataset quality estimation.
+We average samples’ mimic scores in each curated dataset to estimate dataset quality and compare with their corresponding performance gains.
+The results, shown in Figure 4 (right), reveal a positive alignment with a Pearson correlation of 0.958, demonstrating the feasibility to use mimic score as a reliable metric to predict potential model performance gains.
 Sample Analysis.
-We analyze high-value and low-value samples ranked by mimic score.
+We analyze high- and low-value samples ranked by mimic score.
 In Figure 5, we compare token usage in captions.
 We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples.
-Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or incoherent images in the medium-scale dataset (see examples in Figure 1).
+Moreover, mimic score-based filter effectively identifies low-value samples, such as misaligned-caption images in the medium-scale dataset (see examples in Figure 1).
 These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
@@
 Furthermore, we calculate their correlation on random samples.
-We normalize both scores into the range 0 to 1 for clarity The observed positive correlation supports mimic scores in identifying high-quality samples.
-Figure 5: The filtering logic behind Grad-Mimic aligns with handcrafted heuristics, particularly in caption length.
-Low-value samples (left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
-((a))
-((b))
-Figure 6: Mimic score distribution and correlation to CLIP score.
+We normalize both scores into the range 0 to 1 for clarity.
+The observed positive correlation supports mimic scores in identifying high-quality samples (those have high CLIP score).
+Figure 5: The left two figures compare token counts of bottom- and top-ranked samples by mimic score.
+The filtering logic matches handcrafted heuristics, particularly in caption length, where low-value samples have captions that are either too short or over the 77-token limit.
+The right two figures show mimic score distribution and its correlation with CLIP score.
+Table 5: Mimic score can help in predicting whether a sample was used to train the reference model.
 Filtering Strategy Random Selection Mimic Score Selection
@@
 CLIP-ViT L/14 Top 30% 0.150 0.260 0.236 0.382
-DataComp’23 Top-ranked 0.166 0.284 0.271 0.426
-Table 5: Mimic score can help in predicting whether a sample was used to train the reference model.
-Predicting Pretraining Dataset
+DataComp’23 Top-ranked [55] 0.166 0.284 0.271 0.426
+Predicting Pretraining Dataset.
 We explore an interesting application of mimic score: predicting the pretraining dataset.
@@
 As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling.
-We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking final-layer weights without direct access to their filtering steps.
+We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset [55] by simply mimicking final-layer weights without direct access to their filtering steps.
 6 Conclusion
-We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model training.
-We quantify samples by measuring the alignment between gradients and the vector toward pretrained reference weights.
-Using this, we develop Grad-Mimic to automate data selection.
-Empirically, Grad-Mimic effectively prioritizes samples for learning, improving data efficiency and model performance.
-Moreover, Mimic Score aids in creating effective filters and estimating dataset quality.
-Resulting filters outperform human-designed ones and exhibit a strong correlation with dataset quality and performance gains.
-7 Impact Statements
-Efficient data selection is increasingly important for training foundation models.
-We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and resulting model performance.
-We do not foresee explicit negative social impacts arising from Grad-Mimic.
-However, if the reference model used to evaluate sample utility is suboptimal, there remains a risk of refining low-quality datasets (we have an analysis of the impact of reference model quality in Appendix C).
-We believe that careful selection and validation of this reference model, a key area of future research, will minimize this risk and enable the creation of improved datasets.
+We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model weight updates.
+By measuring the alignment of gradients with the vector toward pretrained reference weights, Mimic Score enables precise data assessment.
+Leveraging this scoring metric, we develop Grad-Mimic, a two-stage framework that automates high-quality data selection.
+Theoretically, we study how mimic score-guided training can converge faster than gradient descent.
+Empirically, Grad-Mimic effectively improves data efficiency, accelerating model convergence, and accurately detect mislabeled samples.
+The resulting filters outperform human-designed ones, and their statistics exhibit a strong correlation to training dataset quality and model performance improvements.
 7 Acknowledgments
-We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, John Cooper, Harit Vishwakarma, Dyah Adila, Changho Shin, Satya Sai Srinath Namburi, Zhiqi Gao, Avi Trost, and Gabe Orlanski for their helpful feedback and valuable discussion.
-Appendix A Grad-Mimic Algorithm
-We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights.
-Our algorithm supports various training paradigms, including both supervised learning and self-supervised learning.
-We validate its effectiveness through controlled simulation experiments (Sec.
-5.1) and large-scale training of CLIP models on web-crawled datasets (Sec.
-5.2).
-Furthermore, our experiments encompass a range of training configurations, such as full-parameter fine-tuning, linear probing, and training models from scratch.
-Algorithm 1 Grad-Mimic (Stage 1)
-reference model $\theta_{\text{ref}}$, new model $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature $\tau$
-normalized mimic scores $\overline{m}_{i,t}$, a well-trained model $\theta_{T}$
-Initialize new weights $\theta_{0}$
-for $t\in[T]$ do
-     uniformly sample a batch $\{s_{i}\}_{i=1}^{b}$ from $D$
-     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as a guide
-     for $i\in[b]$ do
-         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradient
-         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright compute mimic score
-     end for
-     $\overline{m}_{:,t}=\frac{e^{m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}$triangleright normalize mimic scores
-     $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright reweight gradients
-end for
-Appendix B Experimental Details
-We provide more details about our training setups and computational resources.
-B.1 Simulation Experiments
-In the experiments of simulating mislabeled samples (Sec.
+We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, Harit Vishwakarma, Dyah Adila, Changho Shin, Satya Sai Srinath Namburi, Zhiqi Gao, Avi Trost, and Gabe Orlanski for their helpful feedback and valuable discussion.
+Appendix A Theoretical Analysis
+We first present a glossary table listing the defined notations and their meanings used in the following analysis.
+Symbol Meaning
+$\mathbb{E}$ Expectation over inherent randomness, such as over gradient noise $\delta$.
+$\mathbb{E}^{(x)}$ Expectation over the dataset
+$\sigma(\cdot)$ The softmax function
+$n$ Dataset size
+$L$ Smoothness parameter on loss function $\ell$
+$\delta$ Level of the gradient noise, $\delta=\mathbb{E}[\|\delta_{i,t}\|^{2}]$
+$\epsilon$ Parameter distance between reference model and the optimal model, $\epsilon=\|\theta_{\text{ref}}-\theta_{*}\|$
+$\tau$ Temperature parameter used in Grad-Mimic updates
+$v_{t}$ Target vector induced by reference model, $\theta_{\text{ref}}-\theta_{t}$
+$-a_{i,t}$ The alignment of the gradient of the loss with the target vector, $g_{i,t}^{\top}v_{t}$
+$\lambda_{t}$ The softmax parameter, $1/(\tau\|v_{t}\|)$
+$\ell(\theta_{t};s_{i})$ The loss of the model with parameters $\theta_{t}$ on a given sample $s_{i}$
+$g_{i,t}$ $\nabla\ell(\theta_{t};s_{i})$
+$\tilde{g}_{i,t}$ $\nabla\ell(\theta_{t};s_{i})+\delta_{i,t}$ where $\delta_{i,t}$ is a noise term
+$\theta^{\text{gd}}_{t+1}$ The learned parameter through a standard GD update
+$\theta^{\text{gm}}_{t+1}$ The learned parameter through a Grad-Mimic update
+We aim to analyze the convergence rates of standard Gradient Descent (GD) and the Grad-Mimic (GM) online reweighting algorithm for an $L$-smooth loss function $\ell$, using noisy gradients with noise level $\delta$.
+Our goal is to compare their convergence to the global minimum $\theta_{*}$, explicitly incorporating the term $\epsilon=\|\theta_{\text{ref}}-\theta_{*}\|$ for Grad-Mimic.
+We start by specifying the update rules for both algorithms, then present three lemmas concerning adaptive algorithms and the softmax function.
+Finally, we discuss how these results apply to Grad-Mimic and quantify the improvement over GD.
+Problem Setup. Consider a (potentially noisy) dataset $D=\{s_{i}\}_{i=1}^{n}$ and an $L$-smooth loss function $\ell(\theta;s)$, defined as:
+$\ell(\theta)=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta;s_{i}).$
+The function $\ell$ is $L$-smooth, and its global minimum is denoted by $\theta_{*}$. We assume access to noisy gradients:
+$\tilde{g}_{i,t}=g_{i,t}+\delta_{i,t},$
+where the gradient noise $\delta_{i,t}$ satisfies:
+$\mathbb{E}[\delta_{i,t}]=0,\quad\mathbb{E}[\|\delta_{i,t}\|^{2}]=\delta.$
+Additionally, we have a reference solution $\theta_{\text{ref}}$ closer to the optimum than the initial weight $\theta_{0}$ in the weight space:
+$\|\theta_{0}-\theta_{*}\|\gg\|\theta_{\text{ref}}-\theta_{*}\|=\epsilon.$
+Standard Gradient Descent. An update rule for GD at iteration $t$ is:
+$\theta^{\text{gd}}_{t+1}=\theta_{t}-\eta\frac{1}{n}\sum_{i=1}^{n}\tilde{g}_{i,t}.$ (4)
+Grad-Mimic’s Online Reweighting. An update rule for Grad-Mimic at iteration $t$ is:
+$\theta^{\text{gm}}_{t+1}=\theta_{t}-\frac{\eta}{\sum_{i=1}^{n}\alpha_{i,t}}\sum_{i=1}^{n}\alpha_{i,t}\tilde{g}_{i,t},$ (5)
+with adaptive weighting that:
+$v_{t}$ $=\theta_{\text{ref}}-\theta_{t},$
+$m_{i,t}$ $=-\tilde{g}_{i,t}^{\top}v_{t}/\|v_{t}\|,$
+$\alpha_{i,t}$ $=e^{m_{i,t}/\tau}.$
+This adaptive weighting leverages the proximity of the reference point $\theta_{\text{ref}}$ to the true minimum $\theta_{*}$, which is expected to effectively mitigating gradient noise.
+When it is convenient, the following notation is used:
+$a_{i,t}$ $=-\tilde{g}_{i,t}^{\top}v_{t},$
+$\lambda$ $=1/(\tau\|v_{t}\|),$
+so that
+$\alpha_{i,t}$ $=\exp(\lambda a_{i,t}).$
+We begin with a lemma about any adaptive weighting procedure, where the $\alpha_{i,t}$ (written as $\alpha_{i}$ unless the timestep needs to be specified) terms are unconstrained.
+Lemma A.1.
+Let
+$R=2\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},a_{i,t})-\eta(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2})$
+Then
+$\|\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}}\|^{2}=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}-\eta R$
+Proof.
+First, begin with a calculation of what $\|\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}}\|^{2}$ is.
+$\|\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}}\|^{2}$ $=\|\theta_{t}-\eta\frac{\sum_{i=1}^{n}\alpha_{i}\tilde{g}_{i,t}}{\sum_{i=1}^{n}\alpha_{i}}-\theta_{\text{ref}}\|^{2}$
+$=\|\theta_{t}-\eta\frac{\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}-\theta_{\text{ref}}\|^{2}$
+$=\|\theta_{t}-\theta_{\text{ref}}\|^{2}-2\eta(\theta_{t}-\theta_{\text{ref}})^{\top}\frac{\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}+\eta^{2}\left\|\frac{\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}\right\|^{2}$
+$=\|\theta_{t}-\theta_{\text{ref}}\|^{2}+2\eta\frac{\mathbb{E}^{(x)}[\alpha_{i}v_{t}^{\top}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}+\eta^{2}\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+Next, using the following identities,
+$\frac{\mathbb{E}^{(x)}[\alpha_{i}v_{t}^{\top}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}$ $=\frac{\mathbb{E}^{(x)}[\alpha_{i}v_{t}^{\top}\tilde{g}_{i,t}]-\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]+\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}$
+$=\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]+\frac{\mathbb{E}^{(x)}[\alpha_{i}v_{t}^{\top}\tilde{g}_{i,t}]-\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]}{\mathbb{E}^{(x)}[\alpha_{i}]}$
+$=\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]+\frac{\text{Cov}^{(x)}(\alpha_{i},v_{t}^{\top}\tilde{g}_{i,t})}{\mathbb{E}^{(x)}[\alpha_{i}]}$
+and
+$\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$ $=\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\mathbb{E}^{(x)}[\alpha_{i}]^{2}\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}+\mathbb{E}^{(x)}[\alpha_{i}]^{2}\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+$=\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}+\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\mathbb{E}^{(x)}[\alpha_{i}]^{2}\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+$=\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}+\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+We can proceed as
+$\|\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}}\|^{2}$ $=\|\theta_{t}-\theta_{\text{ref}}\|^{2}+2\eta\left(\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]+\frac{\text{Cov}^{(x)}(\alpha_{i},v_{t}^{\top}\tilde{g}_{i,t})}{\mathbb{E}^{(x)}[\alpha_{i}]}\right)$
+$+\eta^{2}\left(\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}+\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}\right)$
+$=\|\theta_{t}-\theta_{\text{ref}}\|^{2}+2\eta\mathbb{E}^{(x)}[v_{t}^{\top}\tilde{g}_{i,t}]+\eta^{2}\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}$
+$+2\eta\frac{\text{Cov}^{(x)}(\alpha_{i},v_{t}^{\top}\tilde{g}_{i,t})}{\mathbb{E}^{(x)}[\alpha_{i}]}+\eta^{2}\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}+2\eta\frac{\text{Cov}^{(x)}(\alpha_{i},v_{t}^{\top}\tilde{g}_{i,t})}{\mathbb{E}^{(x)}[\alpha_{i}]}+\eta^{2}\frac{\|\mathbb{E}^{(x)}[\alpha_{i}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\alpha_{i}]\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2}}{\mathbb{E}^{(x)}[\alpha_{i}]^{2}}$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}+2\eta\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},v_{t}^{\top}\tilde{g}_{i,t})+\eta^{2}(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2})$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}-2\eta\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},a_{i,t})+\eta^{2}(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}_{i,t}]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}_{i,t}]\|^{2})$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}-\eta R$
+From Lemma A.1, we want to bound $R$ to hopefully show that $R>0$, indicating that taking an update step with Grad-Mimic will converge faster than GD.
+Lemma A.2.
+Let $f(\lambda)=\frac{\sum_{i=1}^{n}\exp(\lambda v_{i})v_{i}}{\sum_{i=1}^{n}\exp(\lambda v_{i})}$ and $\sigma(\lambda)$ be the vector where $\sigma(\lambda)_{i}=\frac{\exp(\lambda v_{i})}{\sum_{j=1}^{n}\exp(\lambda v_{j})}$.
+Then
+$f(\lambda)-f(0)\geq\frac{1}{\lambda}\|\sigma(\lambda)-\sigma(0)\|^{2}$
+Proof.
+Following the definitions,
+$f(\lambda)-f(0)$ $=\sigma(\lambda)^{\top}v-\sigma(0)^{\top}v$
+$=(\sigma(\lambda)-\sigma(0))^{\top}(v-0)$
+Now, fix a $\lambda$.
+Abusing notation, let $\sigma(v)_{i}=\frac{\exp(\lambda v_{i})}{\sum_{j=1}^{n}\exp(\lambda v_{j})}$, where $\lambda$ is now fixed and the vector $v$ is variable.
+As such, $f(\lambda)-f(0)=(\sigma(v)-\sigma(0))^{\top}(v-0)$.
+Since $\sigma(\cdot)$ is $\lambda$-Lipschitz, it follows that
+$(\sigma(v)-\sigma(0))^{\top}(v-0)\geq\frac{1}{\lambda}\|\sigma(v)-\sigma(0)\|^{2}$
+then we can have
+$f(\lambda)-f(0)\geq\frac{1}{\lambda}\|\sigma(\lambda)-\sigma(0)\|^{2}$
+Lemma A.3.
+The value of $R$ in Grad-Mimic is guaranteed to be positive if
+$\eta<\frac{2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}}{\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2}},$
+or $\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}\leq\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2}$.
+Proof.
+First, recall
+$R=2\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},(\theta_{t}-\theta_{\text{ref}})^{\top}\tilde{g}(\theta_{t};s_{i}))-\eta(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2})$
+Inspecting the first term, as a result of the previous lemma,
+$\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},(\theta_{t}-\theta_{\text{ref}})^{\top}\tilde{g}(\theta_{t};s_{i}))$ $=\text{Cov}^{(x)}(\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]},\tilde{a}_{i,t})$
+$=\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{a}_{i,t}]-\mathbb{E}^{(x)}[\tilde{a}_{i,t}]$
+$=\frac{\sum_{i=1}^{n}\alpha_{i}\tilde{a}_{i,t}}{\sum_{i=1}^{n}\alpha_{i}}-\frac{\sum_{i=1}^{n}\exp(0)\tilde{a}_{i,t}}{\sum_{i=1}^{n}\exp(0)}$
+$=\frac{\sum_{i=1}^{n}\exp(\lambda\tilde{a}_{i,t})\tilde{a}_{i,t}}{\sum_{i=1}^{n}\exp(\lambda\tilde{a}_{i,t})}-\frac{\sum_{i=1}^{n}\exp(0\tilde{a}_{i,t})\tilde{a}_{i,t}}{\sum_{i=1}^{n}\exp(0\tilde{a}_{i,t})}$
+$=f(\lambda)-f(0)$
+$\geq\frac{1}{\lambda}\|\sigma(\lambda)-\sigma(0)\|^{2}$
+Thus,
+$R\geq\frac{2}{\lambda}\|\sigma(\lambda)-\sigma(0)\|^{2}-\eta(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2})$
+Finally, with the definition $\lambda=1/(\tau\|v_{t}\|)$,
+$R\geq 2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}-\eta(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2})$
+A sufficient condition for $R>0$ is therefore given by
+$\eta<\frac{2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}}{\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2}}$
+So far, the only quantity being bounded is the distance to the reference model $\theta_{\text{ref}}$.
+Lastly, we extend to bound convergence to the optimum $\theta_{*}$ and present our Theorem A.4 by using 3 lemma above.
+Theorem A.4.
+Let $\kappa=\max_{i}\|\tilde{g}(\theta_{t};s_{i})\|$. Then
+$\|\theta^{\text{gm}}_{t+1}-\theta_{*}\|^{2}\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}-\eta\left[2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}-4\epsilon\kappa-\eta\kappa^{2}\right]$
+Specifically, if $\eta<\frac{2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}-4\epsilon\kappa}{\kappa^{2}}$ then $\|\theta^{\text{gm}}_{t+1}-\theta_{*}\|\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|$.
+Proof.
+Starting with an application of Lemma A.1,
+$\|\theta^{\text{gm}}_{t+1}-\theta_{*}\|^{2}$ $=\|\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}}\|^{2}+\|\theta_{\text{ref}}-\theta_{*}\|^{2}+2(\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}})^{\top}(\theta_{\text{ref}}-\theta_{*})$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}}\|^{2}+\|\theta_{\text{ref}}-\theta_{*}\|^{2}+2(\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}})^{\top}(\theta_{\text{ref}}-\theta_{*})-\eta R$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}-2(\theta^{\text{gd}}_{t+1}-\theta_{\text{ref}})^{\top}(\theta_{\text{ref}}-\theta_{*})+2(\theta^{\text{gm}}_{t+1}-\theta_{\text{ref}})^{\top}(\theta_{\text{ref}}-\theta_{*})-\eta R$
+$=\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+2(\theta^{\text{gm}}_{t+1}-\theta^{\text{gd}}_{t+1})^{\top}(\theta_{\text{ref}}-\theta_{*})-\eta R$
+$\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+2\epsilon\|\theta^{\text{gd}}_{t+1}-\theta^{\text{gm}}_{t+1}\|-\eta R$
+Thus, we achieve superior performance towards the optimum if
+$R>\frac{2\epsilon\|\theta^{\text{gm}}_{t+1}-\theta^{\text{gd}}_{t+1}\|}{\eta}$
+Using Lemma A.3,
+$\|\theta^{\text{gm}}_{t+1}-\theta_{*}\|^{2}$ $\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+2\epsilon\|\theta^{\text{gd}}_{t+1}-\theta^{\text{gm}}_{t+1}\|-\eta R$
+$\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+2\epsilon\|\theta^{\text{gd}}_{t+1}-\theta^{\text{gm}}_{t+1}\|-2\eta\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}$
+$+\eta^{2}(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2})$
+$\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+4\epsilon\eta\kappa-2\eta\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}$
+$+\eta^{2}(\|\mathbb{E}^{(x)}[\frac{\alpha_{i}}{\mathbb{E}^{(x)}[\alpha_{i}]}\tilde{g}(\theta_{t};s_{i})]\|^{2}-\|\mathbb{E}^{(x)}[\tilde{g}(\theta_{t};s_{i})]\|^{2})$
+$\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}+4\epsilon\eta\kappa-2\eta\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}+\eta^{2}\kappa^{2}$
+where the penultimate inequality uses $\|\theta^{\text{gm}}_{t+1}-\theta^{\text{gd}}_{t+1}\|\leq\|\theta^{\text{gm}}_{t+1}\|+\|\theta^{\text{gd}}_{t+1}\|\leq 2\eta\kappa$.
+We then can find a bound of $\eta$ for $\|\theta^{\text{gm}}_{t+1}-\theta_{*}\|^{2}\leq\|\theta^{\text{gd}}_{t+1}-\theta_{*}\|^{2}$:
+$\eta<\frac{2\tau\|v_{t}\|\cdot\|\sigma(1/\tau\|v_{t}\|)-\sigma(0)\|^{2}-4\epsilon\kappa}{\kappa^{2}}.$
+Theorem A.4 Discussion.
+Our convergence bound reveals conditions under which Grad-Mimic outperforms GD.
+First, the learning rate $\eta$ should be small.
+A large $\eta$ causes the $\kappa^{2}$ term to dominate, leading to large, erroneous steps in Grad-Mimic compared to GD.
+This is particularly evident when the temperature is small, as Grad-Mimic amplifies specific gradients while GD averages conflicting ones, potentially overshooting the optimum if the gradient magnitude $\kappa$ is large.
+Second, $\kappa$ should be small to avoid similar overshooting issues.
+Third, $\epsilon=\|\theta_{\text{ref}}-\theta_{*}\|$ should be small.
+This follows from our intuition: a pre-trained reference model $\theta_{\text{ref}}$ serves as a proxy to the optimum $\theta_{*}$, prioritizing weight movements toward this proxy can accelerate convergence.
+Appendix B Grad-Mimic Algorithms
+Here, we summarize algorithm steps.
+Grad-Mimic proceeds in two stages.
+Stage 1 (Sec.
+4.1) computes mimic scores on the fly and updates model weights through re-weighting.
+Stage 2 (Sec.
+4.2) binarizes mimic scores and aggregates utility assessments to obtain the final estimate.
+Our algorithm supports various training paradigms, including supervised learning and self-supervised learning, as demonstrated in Sec.
+5.1 and Sec.
+5.2.
+Our experiments encompass a range of training configurations, including full-parameter fine-tuning, linear probing, and training models from scratch.
+Algorithm 1 Online Batch Reweighting (Stage 1)
+1:Reference model $\theta_{\text{ref}}$, model being trained $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature parameter $\tau$
+2:Normalized mimic scores $\overline{m}_{i,t}$ for each sample $i$ at step $t$
+3:Initialize model weights $\theta_{0}$
+4:for $t=1$ to $T$ do
+5:     Sample a batch $\{s_{i}\}_{i=1}^{b}$ uniformly from $D$
+6:     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright Weight difference vector
+7:     for $i=1$ to $b$ do
+8:         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright Per-sample gradient
+9:         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright Mimic score
+10:     end for
+11:     $\overline{m}_{:,t}\leftarrow\frac{\exp(m_{:,t}/\tau)}{\sum_{j=1}^{b}\exp(m_{j,t}/\tau)}$triangleright Normalize mimic scores
+12:     $\theta_{t+1}\leftarrow\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright Update weights with reweighted gradients
+13:end for
+Algorithm 2 Offline Sample Selection (Stage 2)
+[…]
+2:Aggregated retain probabilities $p_{i}$ for each sample $i$, retained subset $\mathcal{S}$
+3:Initialize binary decision matrix $B\leftarrow\mathbf{0}^{n\times T}$
+4:for $t=1$ to $T$ do
+5:     $m\leftarrow\overline{M}_{:,t}$triangleright Normalized mimic scores at step $t$
+[…]
+7:         $\tau\leftarrow 1/b$triangleright Threshold based on batch size
+8:         $B_{:,t}\leftarrow\mathbf{1}[m>\tau]$
+[…]
+10:         Fit a two-cluster $k$-means model on $m$triangleright 1D clustering
+11:         $B_{:,t}\leftarrow$ cluster assignments
+[…]
+13:         $\tau\leftarrow k$-th percentile of $m$
+14:         $B_{:,t}\leftarrow\mathbf{1}[m\geq\tau]$
+15:     end if
+16:end for
+17:Train a Snorkel LabelModel $\mathcal{LM}$ on $B$
+18:$P\leftarrow\mathcal{LM}.\text{predict\_proba}(B)$triangleright Retain/discard probabilities, $P\in\mathbb{R}^{N\times 2}$
+19:$p_{i}\leftarrow P_{i,1}$ for each sample $i$triangleright Extract retain probabilities
+20:$\mathcal{S}\leftarrow\{i\mid p_{i}>0.5\}$triangleright Default threshold $0.5$
+21:return $\{p_{i}\}_{i=1}^{n}$, $\mathcal{S}$
+Appendix C Experimental Details
+We shift the focus from theoretical analysis and algorithm steps to empirical results. We first detail the training configurations and computational resources used in our experiments.
+Mislabeled Sample Detection.
+For the experiments of simulating mislabeled samples (Sec.
 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [56].
-Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs.
-We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper.
-Results for other temperature values are presented in Table 7.
-To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets but with different random seeds for weight initialization.
-These experiments were performed on a Nvidia Tesla A100.
-B.2 DataComp Experiments
-We adopt the training setup from DataComp [49, 17] and train CLIP models from scratch using contrastive objective over image-caption pairs.
-Each CLIP model is trained for 5 epochs with a batch size of 4096.
+Training is conducted with a batch size of 32, a learning rate of 1e-4, and the AdamW optimizer.
+To ensure convergence, we set the training steps to 10 for the DTD and Flowers102 datasets and to 5 for the remaining datasets.
+We evaluate Grad-Mimic across various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report performance at $\tau=0.5$ in the main paper (Table 1).
+Results for other temperature values are provided in Table 9.
+To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets, using different random seeds for weight initialization.
+All experiments are conducted on an NVIDIA Tesla A100.
+Large-scale Data Curation.
+For the experiments of multimodal data selection (Sec.
+5.2), we follow the training protocol used in DataComp [54, 21] and train CLIP models from scratch using contrastive objective on image-caption pairs.
+Each model is trained for 5 epochs with a batch size of 4096.
 The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset.
-These experiments run on 8 Nvidia Tesla A100s.
-We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders.
-Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [17] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5).
-We report results at $\tau=0.5$ and $\tau=0.05$ in the main body in Table 3, but full results are presented in Table 6.
+We analyze different components of the reference model’s weights, specifically the final MLP layers in the image and text encoders.
+We evaluate Grad-Mimic’s effectiveness using temperature values $\tau$ of 0.03, 0.05, 0.07, 0.3, and 0.5.
+Performance is assessed across 38 diverse downstream tasks [21].
+Results for $\tau=0.5$ and $\tau=0.05$ are reported in the main paper (Table 3), and comprehensive results are provided in Table 8.
+These experiments are conducted on 8 NVIDIA Tesla A100 GPUs.
+Appendix D Comparing to Other Model-based Approaches
+We compare Grad-Mimic with other model-based techniques, specifically reference model-based approach and influence function-based approach.
+Table 6: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing GPU memory usage and substantial computational overhead (see efficiency advantages are in Appendix E).
+DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
+Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
+Mini-Batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
+Rho-Loss 54.10 52.39 48.46 41.55 35.71 30.02 97.10 96.83 96.49 88.53 87.79 86.86 94.07 93.83 93.82 76.82 76.05 75.93 75.36 73.77 71.93
+Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
+Table 7: Grad-Mimic outperforms other reference model-based methods even when using weaker or distilled reference model as proxy.
+CIFAR10 CIFAR100
+Noise Level 0.5 0.5
+Rho-Loss 93.83 76.05
+Using full training set Grad-Mimic 93.92 76.82
+Rho-Loss 93.26 75.11
+Using hold-out training set Grad-Mimic 93.38 75.81
+Figure 6: Grad-Mimic outperforms approximated influence function-based methods: Using reference model weights as selection guide offer advantages on their effectiveness and greater accessibility.
+Reference Model-Based Approaches.
+We compare Grad-Mimic to Rho-Loss [14], a representative method that prioritizes samples based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on a pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$.
+For a fair comparison, we simulate pretrained reference models using two approaches: (i) training on a noise-free hold-out dataset as used in Rho-Loss [14], and (ii) training on the entire training dataset, consistent with our setup in the main paper.
+We adopt our mislabeled sample detection setup and evaluate the model performance performance on the testing datasets.
+We first demonstrate model performance using reference model that is trained on entire training dataset.
+As shown in Table 6, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels.
+Next, we benchmark both methods on CIFAR10 and CIFAR100 with 50% label noise, using a reference model trained on a noise-free hold-out dataset.
+Such setup simulates more realistic scenarios where the reference model, acting as a distilled or weaker approximation of ideal weights, is used.
+The results, presented in Table 7, again show that Grad-Mimic achieves higher test accuracy than Rho-Loss.
+Moreover, Grad-Mimic offers other computational advantages.
+Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes to compute excess loss, Grad-Mimic scores sample utility directly from discrepancies in model weights—a simpler and more efficient operation.
+Overall, these results are showing that using only a subset of weights, such as those in the last layer, is sufficient for Grad-Mimic to ensure high performance, further reducing memory usage and improving computational efficiency.
+Influence Function-based Approaches.
+Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [57, 58, 59], but their computational cost often limits practical use.
+Recent studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a noisy-free validation dataset [19, 20, 18].
+We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from a noise-free validation dataset.
+We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from per class (1, 5, 10, 15, and 20).
+Results in Figure 6 show that while larger validation sets enhance data efficiency by enabling models to focus on influential samples, Grad-Mimic, which leverages pretrained reference weights and their induced direction, consistently outperforms influence approximations based on validation datasets.
+Moreover, relying on labeled validation sets introduces dataset dependencies, incur computational overhead from validation dataset’s gradient computations, and poses challenges in ensuring their quality, as noted in the limitations of GREATS [18].
+For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2,000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
+Appendix E Efficiency Advantages
+Figure 7: Grad-Mimic enjoys lower runtime per batch and smaller memory usage than other baselines.
+Figure 8: Grad-Mimic leads to faster convergence and improved data utilization.
+We provide four distinct lines of evidence for the efficiency gains our technique. They are (i) computational efficiency, (ii) memory usage, (iii) learning efficiency, and (iv) data utilization.
+Computational Efficiency and Memory Usage.
+First, we benchmark per-batch training runtime and peak GPU memory consumption across considered baselines, as presented in Figure 7.
+Grad-Mimic achieves computational runtime and memory usage comparable to GraNd, relying solely on per-sample training gradients.
+In contrast, methods like AGRA and approximated influence function-based approaches incur higher computational costs due to additional gradient computations from either another training batch or a noise-free validation dataset.
+This creates a 2.6 times overhead.
+Similarly, Grad-Mimic outperforms Rho-Loss in efficiency, as Rho-Loss requires loading a full reference model and performing additional forward passes—substantially increasing compute time.
+Regarding memory usage, Rho-Loss can exceed all these methods if the reference model size scales significantly.
+Learning Efficiency.
+Second, we evaluate learning efficiency by analyzing convergence curves.
+As shown in Figure 8 (left), experiments on CIFAR100 with 50% label noise demonstrate that Grad-Mimic converges faster than all compared methods, achieving higher model performance within a fixed number of training epochs.
+Additionally, Figure 4 (left) illustrates that, on the DataComp medium-scale dataset (over 100 million samples), Grad-Mimic, aided by a reference model, converges faster than standard training, reducing training steps by 20.7%.
+Data Utilization.
+Finally, we introduce a new experiment that measures data efficiency without adding label noise.
+Instead of simulating mislabeled samples for denoising, we rank clean training samples by their mimic scores and build subsets from the top-kentries.
+We use a well-trained ViT-B/16 as our reference model and perform Grad-Mimic on clean CIFAR100 set.
+We evaluate the model performance trained on refined subsets.
+Results, presented in Figure 8, show that Grad-Mimic’s selected subsets consistently outperform randomly drawn subsets of the same size.
+Remarkably, with both ViT-B/16 and the larger ViT-L/14, Grad-Mimic matches the full-dataset accuracy while reducing the training data by 20%.
+This highlights a direct reduction in data-related computational cost.
+These findings—lower runtime per batch, reduced memory usage, faster convergence, and improved data utilization—demonstrate that Grad-Mimic achieves significant efficiency gains, fulfilling our motivation.
+Table 8: Full Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance across all temperature settings.
 Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
@@
 0.05 0.169 0.151 0.262 0.216 0.258
-Table 6: Full Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance across all temperature settings.
-Appendix C Ablation Studies
+Appendix F Ablation Studies
+Table 9: Stage 1 Results in Simulation Experiment using Various Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
 Temperature $\tau$ DTD Flower102 Oxford-IIIT Pet CIFAR10 CIFAR100
@@
 Grad-Mimic 0.6 56.0 46.4 89.3 94.2 77.3
-Table 7: Stage 1 Results in Simulation Experiment using Various Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
-The Choice of Temperature.
-We test Grad-Mimic using different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [24]).
+Choice of Temperature.
+We run Grad-Mimic under different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [29]).
 The results in the simulation experiments are presented in Table 9.
 In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration.
-Grad-Mimic outperforms the baseline methods in all temperature settings.
+Grad-Mimic outperforms both baseline methods at every temperature.
 Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
-DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
-Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
-Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
-Mini-Batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
-Rho-Loss 54.10 52.39 48.46 41.55 35.71 30.02 97.10 96.83 96.49 88.53 87.79 86.86 94.07 93.83 93.82 76.82 76.05 75.93 75.36 73.77 71.93
-Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
-Table 8: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing memory usage and computational overhead.
-Comparison to Reference Model-Based Approaches.
-We compare Grad-Mimic to Rho-Loss [15], a method that prioritizes samples for learning based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on the pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$.
-We follow our simulation experiment setup and evaluate the performance of fine-tuned models on the testing datasets.
-As shown in Table 8, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels.
-Furthermore, Grad-Mimic offers efficiency advantages.
-Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes, Grad-Mimic computes mimic scores directly from the reference model’s weights.
-In our experiments, we demonstrate that even using only a subset of the weights (e.g., the last-layer weights) is sufficient, further reducing memory usage and improving computational efficiency.
-Figure 7: Grad-Mimic can outperform influence function-based methods: Using model weights as a guide offer advantages on their effectiveness and greater accessibility.
-Comparison to Influence Function-based Approaches.
-Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [52, 53], but their computational cost often limits practical use.
-Several studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a validation dataset [54, 55].
-We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from validation dataset.
-We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from each class (1, 5, 10, 15, and 20).
-Results in Figure 7 show that while larger validation sets improve data efficiency by enabling model to focus on the influential samples, Grad-Mimic—using pretrained reference weights and their induced direction—consistently outperforms validation dataset-based influence approximations.
-Moreover, relying on a labeled validation set risk introducing an unwanted dependency.
-For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
-Figure 8: Grad-Mimic can be robust to the quality of reference model.
+Table 10: Last-layer weights yields the best performance across all datasets.
+DTD Oxford-IIIT Pet CIFAR10 CIFAR100
+Layer 7 0.575 0.755 0.907 0.752
+Layer 8 0.583 0.755 0.908 0.753
+Layer 9 0.586 0.740 0.911 0.749
+Layer 10 0.582 0.751 0.902 0.753
+Layer 11 (last-layer weights) 0.595 0.758 0.912 0.755
+Which layer should be mimicked?
+Next, we study the impact of the layer weights to mimic.
+We vary the depth of the reference model weights used to obtain the target vector and evaluated the resulting model performance.
+Results, presented in the Table 10, indicate that mimicking the last-layer weights yields the highest accuracy in most cases.
+While the optimal choice may depend on architecture or task, these findings suggest that later representations generally offer the richest information and are better choices for estimating sample utility.
 Impact of Reference Model Quality.
 Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available.
-To simulate such scenarios, we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights.
+To simulate such scenarios, we degrade the reference model in two controlled ways:
+  • Weight corruption: we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights.
 By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
-We present this analysis in Figure 8.
-As expected, increasing noise variance degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic.
-However, we observe that even under significant noise, models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself.
-This highlights that while a high-quality reference model is ideal, even a weaker proxy can provide valuable guidance and improve training outcomes over traditional baselines.
+  • Data-limited training: we train the reference model on progressively smaller fractions of its original dataset (shrinking from 1.3B samples to 3.5M samples).
+Fewer training samples will produce a weaker (or distilled) reference model.
+We assess Grad-Mimic in both the mislabeled-sample and multimodal settings, applying each degradation method separately.
+We present these two analysis in Figure 9 and Table 11.
+As expected, increasing noise variance or using less training data degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic.
+However, we observe that even under significant noise (heavy degradation), models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself (see Figure 9).
+This highlights that while a high-quality reference model is ideal, a weaker proxy can still provide valuable guidance and improve training over traditional baselines.
+Table 11: Grad-Mimic achieves performance gains even when using a weaker reference model.
+Scale # of samples to train $\theta_{\text{ref}}$ Training Method ImageNet
+ImageNet
+dist. shifts
+VTAB Retrieval
+Average over
+38 datasets ($\uparrow$)
+Using 1.3B samples Vanilla Training 0.026 0.035 0.139 0.114 0.131
+Using 1.3B samples Grad-Mimic 0.026 0.037 0.162 0.118 0.146
+Small Using 3.5M samples Grad-Mimic 0.026 0.035 0.158 0.115 0.140
+Figure 9: Grad-Mimic can be robust to the quality of reference model.
+Appendix G Broader Impacts and Limitations
+Efficient data selection is a crucial prerequisite for training multimodal models.
+We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and downstream model performance.
+We do not foresee any direct negative social consequences from the technique itself.
+However, if the reference model used to evaluate sample utility is poorly calibrated or biased, there remains a risk of refining low-quality datasets.
+Appendix F investigates such scenario.
+It shows that Grad-Mimic remains competitive even when the reference model is heavily degraded, yet some performance loss is unavoidable.
+We believe that practitioners can easily mitigate this risk by validating the reference model on a small, trusted benchmark before data scoring.
+This check is easier than the requirements of many recent data-selection methods, which rely on noise-free validation sets.
+By decoupling data scoring from such hard-to-obtain resources, Grad-Mimic offers greater flexibility and control.